### v1

# WKB structure in a scalar model of flat bands

Abstract.
We consider a family of periodic scalar operators for which one can define flat bands in the sense of Floquet–Bloch theory. One puzzling question originating in recent physics literature is a quantisation rule for the values of parameters at which these flat bands occur. We present a general theorem about the structure of solutions to the corresponding equation and a heuristic argument explaining their WKB structure in a specific case. That structure also explains the quantisation condition – both the WKB structure and that rule are confirmed by numerical experiments. Finally, we consider a simplified model in which separation of variables allows the use of complex WKB methods.
1. Introduction
Topologically nontrivial flat bands (in the sense of Bloch–Floquet theory) are of interest in condensed matter physics as their presence has remarkable experimental consequences – see [BiMa11], [TKV19] for theoretical references and [C*18] for experimental results in the context of twisted bilayer graphene.
This paper is concerned with flat bands and the structure of eigenstates for a class of periodic operators on $\mathbb{R}^{2}$. They are of the form
$P(\alpha,k)=(2D_{\bar{z}}+k)^{2}-\alpha^{2}V(z),\ \ 2D_{\bar{z}}=\tfrac{1}{i}(\partial_{x_{1}}+i\partial_{x_{2}}),$ (1.1)
where $x=(x_{1},x_{2})$, $\alpha,k\in\mathbb{C}$, and $V$ is a real analytic, $\Lambda$-periodic potential satisfying additional symmetry assumptions – see (2.5). The study of $P(\alpha,k)$ is equivalent to the study of a system
[…]
Flat bands (see (1.5) below) occur at $\alpha$ if
$\operatorname{Spec}_{L^{2}(\mathbb{C}/\Lambda)}D_{\rm{S}}(\alpha)=\mathbb{C}.$ (1.3)
In addition, the symmetries (2.5) imply the existence of non-trivial solutions to
$D_{\rm{S}}(\alpha)v(\alpha)=0,$ (1.4)
for all $\alpha$. They are called (symmetry) protected states as their existence is due to symmetries of the operator. We are interested in the distribution of $\alpha$’s for which (1.3) holds and in the closely related WKB structure (as $\alpha\to\infty$) of the protected states.
Figure 1. The sets $\mathcal{A}$ (red) and $\mathcal{B}$} ( blue) of Theorem 1 for the scalar model based on the Bistritzer–MacDonald potential (2.3). The multiple circles indicate multiplicity (in the sense of Theorem 1, $m=1,2,3$ all occur). A movie showing the evolution of the bands $k\mapsto E_{j}(\alpha,k)$, $j=1,2$ (over the real axis in $k$) as $\alpha$ varies can be see at https://math.berkeley.edu/~zworski/bands_scalar.mp4. It illustrates the behaviour of bands at real elements of $\mathcal{A}\cup\mathcal{B}$.
Because the properties of $D_{S}(\alpha)$ in (1.2) are described using a scalar operator (1.1), we refer to this model as a scalar model of magic angles/flat bands. It was introduced in [GaZw24] with the hope of having a simpler setting for seeing the distribution of magic angles. That model is derived from the the chiral model of twisted bilayer graphene (TBG) [TKV19], [Be*22], [BHZ24], see §2.1.
The self-adjoint Hamiltonian associated to $D_{\rm{S}}(\alpha)$ is given by
$H(\alpha):=\begin{pmatrix}0&D_{\rm{S}}(\alpha)\\ D_{\rm{S}}(\alpha)^{*}&0\end{pmatrix}$,
and the bands are the eigenvalues of $e^{-i\langle z,k\rangle}H(\alpha)e^{i\langle z,k\rangle}:H^{1}(\mathbb{C}/\Lambda;\mathbb{C}^{4})\to L^{2}(\mathbb{C}/\Lambda,\mathbb{C}^{4})$, $z,k\in\mathbb{C}$, considered as functions of $k$:
$\cdots\leq-E_{j}(\alpha,k)\leq\cdots\leq-E_{1}(\alpha,k)\leq 0\leq E_{1}(\alpha,k)\leq\cdots\leq E_{j}(\alpha,k)\leq\cdots$ (1.5)
(For the general theory of Bloch–Floquet band spectrum see [TaZw23, §5.3] and for an introduction in this context [Zw24, §3.1,3.2].) A flat band occurs when $E_{1}(\alpha,k)\equiv 0$ and that is equivalent to (1.3).
Local solutions to $D_{\rm{S}}(\alpha)u=0$ (or equivalently solution to $P(\alpha,0)u_{0}=0$) are given by local holomorphic sections of a rank-2 holomorphic (topologically trivial) vector bundle over $\mathbb{C}/\Lambda$. We denote it by $\mathscr{E}(\alpha)$ and it is defined in §3.6. The first theorem describes the connection between the structure of this vector bundle and the behaviour of bands:
Theorem 1.
Suppose that $E_{j}(\alpha,k)$ are defined in (1.5) for $D_{\rm{S}}(\alpha)$ with $V$ satisfying (2.5). There exist discrete sets $\mathcal{A},\mathcal{B}\subset\mathbb{C}$ such that we have the following trichotomy:
$\alpha\notin\mathcal{A}\cup\mathcal{B}\Longleftrightarrow\mathscr{E}(\alpha)\text{ is indecomposable }\Longleftrightarrow E_{\pm 1}(\alpha,k)\simeq\pm c|k|^{2}$;
$\alpha\in\mathcal{B}\Longleftrightarrow\mathscr{E}(\alpha)=L_{0}\oplus L_{1},L_{j}\text{ trivial}\Longleftrightarrow E_{\pm j}(\alpha,k)\simeq\pm c_{j}|k|,\ j=1,2$, $E_{3}(\alpha,k)>0$;
$\alpha\in\mathcal{A}\Longleftrightarrow\mathscr{E}(\alpha)=L\oplus L^{*},$ deg L = m $\Longleftrightarrow E_{j}(\alpha,k)\equiv 0,\ |j|\leq m$, $E_{m+1}(\alpha,k)>0$,
where $L_{j}$ and $L$ are line bundles.
We refer to $m$ in the third statement of the theorem as the multiplicity of a magic $\alpha\in\mathcal{A}$. Although the model we consider does not seem to be physically relevant, Li–Yang [LiYa] proved that the same trichotomy appears in the case of twisted multiple layers of graphene – see [Ya23] for a mathematical presentation and references to the physics literature. The fact that double Dirac points can occur at a discrete set of angles was not present in the physics literature. The computations in [LiYa] suggest that Dirac points should occur when the angle of twisting between two bilayer graphene wafers if close to 0.44.
1.1. Quantisation condition for real magic $\alpha$’s
One of the most striking observations made in [TKV19] was a quantisation rule for real magic alphas in the chiral model (the definition is the same as (1.3) but with the operator (2.1)) with the Bistritzer–MacDonald potential (2.2): if $\alpha_{1}<\alpha_{2}<\cdots\alpha_{j}<\cdots$ is the sequence of positive $\alpha$’s for which (2.6) holds, then
$\alpha_{j+1}-\alpha_{j}=\gamma+o(1),\ \ j\to+\infty,\ \ \gamma\simeq\tfrac{3}{2}.$ (1.6)
The more accurate computations made in [Be*22] suggest that $\gamma\simeq 1.515$. In the scalar model (2.4) with $V(z)=U_{\rm{BM}}(z)U_{\rm{BM}}(-z)$ where $U_{\rm{BM}}$ is given by (2.2) we numerically observe the following rule for real elements of $\mathcal{A}$: real magic $\alpha$’s are double (in the sense of Theorem 1), $\alpha_{2j-1}=\alpha_{2j}$, $j=1,2,\cdots$,
$\alpha_{j+1}-\alpha_{j-1}=2\gamma+o(1),\ \ j\to+\infty,$ (1.7)
where $\gamma$ is the same as in (1.6). This ”doubling” phenomenon in the passage from chiral to scalar models occurs for other potentials as well.
In this paper we discuss the connection between the WKB structure of solutions to (1.4) and the distribution of $\alpha$’s in the scalar model. In §4.1 we present the mechanism for zero creation in protected states $v(\alpha)$ for the Bistritzer–MacDonald potential. That is related to the distribution of magic alphas – see §3.3. In §4.2 we discuss the potential
$U(z):=i\overline{U_{\rm{BM}}(z)}^{2},\ \ \ V(z):=U(z)U(-z),$ (1.8)
for which the structure of the equation (1.1) simplifies significantly and for which we have (1.6) with $\gamma=\frac{1}{4}$. That is presented in the next theorem though we stress that we do not provide a full proof in this paper. At this point we are only able to provide a heuristic WKB argument supported by numerical evidence. Hence the next theorem is only a theorem in the “physical” sense.
Theorem 2.
The potential $V$ is given in (1.8) satisfies (2.5) and hence defines a set $\mathcal{A}$ of Theorem 1. If $\mathcal{A}\cap\mathbb{R}_{+}=\{\alpha_{j}\}_{j=1}^{\infty}$, $\alpha_{j+1}\geq\alpha_{j}$ then each $\alpha_{2j-1}=\alpha_{2j}$, $j=1,2,\cdots$ (that is, each real $\alpha$ has multiplicity 22) and
$\alpha_{k}-\alpha_{k-2}=\tfrac{1}{4}+o(1),\ \ \ k\to\infty.$ (1.9)
Numerically computed $\alpha_{j}$’s for the scalar model based on (1.8) are
$k$ $\alpha_{k}$ $\alpha_{k}-\alpha_{k-2}$
1 0.1395
3 0.3803 0.2407
5 0.6281 0.2478
7 0.8772 0.2490
9 1.1267 0.2494
11 1.3764 0.2496
13 1.6262 0.2497
15 1.8760 0.2498
17 2.1259 0.2498
19 2.3758 0.2498
This shows that the heuristic argument for (1.9) is confirmed numerically. The magic alphas for the scalar model and the potential $V$ in (1.8) are shown in Figure 2. They have the same structure (on a different scale) as the magic alphas for the Bistritzer–MacDonald potential shows (in red) in Figure 1. We also remark that for the chiral model with the potential $U$ in (2.1) we have an analogue of (1.6) and (1.7) with $\gamma=\frac{1}{8}$: the spacing is halved and the multiplicities are simple.
Figure 2. The sets $\mathcal{A}$} for the scalar model with the potential $V$ given by (1.8). The multiple circles indicate multiplicity (in the sense of Theorem 1, $m=1,2,3$ all occur). Except for different scaling the structure of the set is the same as for $\mathcal{A}$ shown in red in Figure 1.
1.2. Nonvanishing potentials
We can study properties of solutions to $P(\alpha,k)u=0$ for more general potentials. In that case, we do not expect to have $\alpha$’s for which (1.3) holds but we could ask for which $\alpha$’a and $k$ we have a non-trivial kernel of $P(\alpha,k)$. One case to consider is that of
$V=W^{2}\neq 0,$ (1.10)
as in that case the characteristic variety (that is the set of $(x,\xi)\in T^{*}\mathbb{C}/\Lambda$ such that $q(x,\xi):=(2\bar{\zeta})^{2}-V(z)=0$, $\zeta=\frac{1}{2}(\xi_{1}-i\xi_{2})$, $z=x_{1}+ix_{2}$) is a union of two disjoint tori. Those tori are not Lagrangian unless $\operatorname{Im}\partial_{z}W\equiv 0$.
Assuming (1.10) and that
$W_{0}:=|\mathbb{C}/\Lambda|^{-1}\int_{\mathbb{C}/\Lambda}W(z)dm(z)\neq 0,$ (1.11)
we have the following naïve WKB argument for the existence of non-trivial kernel of $P(\alpha,k)$: we construct two global phase functions by solving
$2\partial_{\bar{z}}\varphi+k=\pm\alpha W(z),\ \ \varphi(z)=\pm(\langle\alpha W_{0},z\rangle+\alpha\psi)-\langle k,z\rangle,\ \ \psi(z+\gamma)=\psi(z).$
(Here $\langle k,z\rangle=\operatorname{Re}(k\bar{z})$.) A naïve quantisation condition is the given by
$\forall\,\gamma\in\Lambda\ \varphi(z+\gamma)\equiv\varphi(z)\!\!\!\!\mod\!2\pi\mathbb{Z}\ {\Longleftrightarrow}\ \alpha\in W_{0}^{-1}(\Lambda^{*}\pm k).$ (1.12)
It is not clear under what assumptions this gives an approximate values of $k$ and $\alpha$ for which we have a nontrivial solution to $P(\alpha,k)u=0$. It certainly does not hold in general – see Figure 3. We remark that for non-self-adjoint operators for which the characteristic variety is a single torus, close in some sense to a Lagrangian torus, Melin and Sjöstrand [MeSj03] obtained Bohr–Sommerfeld quantisation rules for the spectrum. Their methods do not seem to be applicable to our case.
Figure 3. The bottom panels show sets of $\alpha$’s (circles; coloured depending on the momentum variable of $y$) for which $(D_{x}+iD_{y})^{2}-\alpha^{2}W(x)$, $(x,y)\in\mathbb{R}^{2}/2\pi\mathbb{Z}^{2}$ has a nontrival kernel. The top panels show curves on which $\operatorname{Im}(W(\gamma(t))\gamma^{\prime}(t))=0$ (Stokes lines). The first $W$ satisfies (5.1) but does not have Stokes loops (1.14); on the second $W$ has Stokes loops. The quantisation rule (1.12) (shown by black dots) does not apply on the top and is very accurate on the bottom. For an animated version see https://math.berkeley.edu/~zworski/stokes.mp4.
In view of these complications we consider a yet simpler model in which one can use separation of variables and the complex WKB. For that we take $\Lambda:=\mathbb{Z}^{2}$ and $W(x,y)=W(x)$, $(x,y)\in\mathbb{R}^{2}$. Taking the Fourier modes in $y$ we obtain a family of operators depending $\zeta=2\pi mi+k$, where $m\in\mathbb{Z}$ is a Fourier mode in $y$ and $k$ is in the fundamental domain of the dual lattice $2\pi\mathbb{Z}^{2}$:
$P_{\zeta}(\alpha):=(D_{x}+\zeta)^{2}-\alpha^{2}W(x)^{2}.$ (1.13)
We assume that $W(x)$ is an entire function satisfying $W(x+1)=W(x)$. In that case methods of complex WKB can be applied but they normally require precise properties of a given $W(x)$, typically a trigonometric polynomial – see for instance the work of Galtsev–Shafarevich [GaSh] and references given there. For a brief account of complex WKB method and additional references we suggest [Sj19, Chapter 7]. Here we present a general result which gives a condition under which the quantisation rule (1.12) holds:
Theorem 3.
Suppose that $W(x)$ in an entire function satisfying $W(x+1)=W(x)$. If there exists a closed smooth curve ${\gamma}:[0,1]\to\mathbb{C}/\mathbb{Z}$, $\gamma^{\prime}(t)\neq 0$, such that
$\operatorname{Im}{\gamma}^{\prime}(t)W({\gamma}(t))=0,\ \ W(\gamma(t))\neq 0,\ \ t\in[0,1],$ (1.14)
then for $|\alpha|\gg 1$ and $|\operatorname{Re}\zeta|\leq\pi$,
$\ker((D_{x}+\zeta)^{2}-\alpha^{2}W(x))\neq\{0\}\ \Longleftrightarrow\ \exists\,n\in\mathbb{Z}\ \alpha=W_{0}^{-1}(2\pi n\pm\zeta)+\mathcal{O}(|\alpha|^{-1}).$
The restriction on $|\operatorname{Re}\zeta|$ is not essential as for $\ell\in\mathbb{Z}$, $P_{\zeta+2\pi\ell}=e^{-2\pi i\ell x}P_{\zeta}e^{2\pi i\ell x}$. The proof shares many features with standard complex WKB methods but we are asking different questions and allow for a very large class of potentials. A more precise version is given in Theorem 4. An illustration of the result is given in Figure 3. One reason for presenting this case, much simpler than that considered in the heuristic Theorem 2, is to indicate the potential complications for operators involving the potentials $U_{\rm{BM}}$ and $i\overline{U_{\rm{BM}}}^{2}$ for which flat bands are present.
A condition on $W$ which guarantees the existence of $\gamma$ satisfying (1.14) (we call such $\gamma$’s Stokes loops) is given in Proposition 17: $\int_{0}^{1}W(x)dx=W_{0}\neq 0$, and
$W(z)\neq 0,\ \ |\operatorname{Im}(e^{-i\arg W_{0}}W(z))|\leq\operatorname{Re}(e^{-i\arg W_{0}}W(z))\ \ \ \text{ for $|\operatorname{Im}z|\leq 1$.}$ (1.15)
For instance, if $f(z)$ is entire and $f(z+1)=f(z)$ then $W(z)=M+f(z)$ will satisfy (1.15) for $M\geq\sqrt{2}\max_{|\operatorname{Im}z|\leq 1}|f(z)|$.
Notation. We write $f=\mathcal{O}_{\rho}(g)_{H}$ for $\|f\|_{H}\leq C_{\rho}g$, that is the constant depends on $\rho$. When the context is clear either $\rho$ or $H$ may be dropped. We denote by $\mathscr{O}(U)$ holomorphic functions on an open set $U\subset C^{d}$. If $z_{0}\in U$ where $U$ is a topological space we write ${\rm neigh}_{U}(z_{0})$ for an open neighbourhood of $z_{0}$ in $U$.
Acknowledgements. We would like to thank Zhongkai Tao for help with the material in §3.6, John Lott for the reference [Da07] and Mengxuan Yang for comments on the first version of this paper. The authors were partially supported by the Simons Targeted Grant Award No. 896630 “Moiré Materials Magic”, while the first named author was partially supported by NSF CAREER grant DMS-1749858.
2. Review of the scalar model
In this section we review the properties of the scalar model, starting with the chiral model from which it is derived.
2.1. Chiral model
Here we review some of its properties. It is derived from the the chiral model of twisted bilayer graphene (TBG) [TKV19], [Be*22], [BHZ24]:
$\begin{gathered}Q(\alpha,k):=D(\alpha)+k,\ \ D(\alpha):=\begin{pmatrix}2D_{\bar{z}}&\alpha U(z)\\ \alpha U(-z)&2D_{\bar{z}}\end{pmatrix},\ \ \ \Omega=\mathbb{C},\\ 2D_{\bar{z}}=\tfrac{1}{i}(\partial_{{x_{1}}}+i\partial_{x_{2}}),\ \ z={x_{1}}+ix_{2}\in\mathbb{C},\end{gathered}$ (2.1)
where $U$ satisfies
$\begin{gathered}U(z+\gamma)=e^{i\langle\gamma,K\rangle}U(z),\ \ U(\omega z)=\omega U(z),\ \ \overline{U(\bar{z})}=-U(-z),\ \ \ \omega=e^{2\pi i/3},\ \\ \gamma\in\Lambda:=\omega\mathbb{Z}\oplus\mathbb{Z},\ \ K=\tfrac{4}{3}\pi,\ \ \langle z,w\rangle:=\operatorname{Re}(z\bar{w}).\end{gathered}$ (2.2)
The constant $K$ is determined up to sign by
$\omega K\equiv K\not\equiv 0\!\!\!\mod\Lambda^{*},\ \ \ \Lambda^{*}:=\frac{4\pi i}{\sqrt{3}}\Lambda,$
that it is a high symmetry point with respect to the dual lattice. Although the chiral model is not considered accurate for large $\alpha$ (a dimensionless constant proportional the reciprocal of the angle of twisting), it remains popular in physics and its mathematical properties remain puzzling – see [Zw24].
The principal example of $U$ is given by the Bistritzer–MacDonald potential
$U_{\rm{BM}}(z)=-\tfrac{4}{3}\pi i\sum_{\ell=0}^{2}\omega^{\ell}e^{i\langle z,\omega^{\ell}K\rangle}.$ (2.3)
It is, up to a real multiplicative constant, the unique potential satisfying (2.2) and having the lowest Fourier modes.
The scalar model is defined as follows:
$P(\alpha,k):=(2D_{\bar{z}}+k)^{2}-\alpha^{2}V(z),\ \ \ V(z):=U(z)U(-z).$ (2.4)
It is derived from the chiral model by neglecting lower order terms, as $\alpha\to\infty$, in the principally scalar reduction:
[…]
We refer to [HiZw25, §2.2] for the discussion of this reduction and a semiclassical explanation why $\alpha R_{k}(\alpha)$ is a lower order term.
The potential $V$ in (2.4) inherits symmetries from (2.2) but we can formulate them independently as
$V(z)=V(-z),\ \ V(z+\gamma)=V(z),\ \ V(\omega z)=\omega^{2}V(z),\ \ \overline{V(\bar{z})}=V(z).$ (2.5)
Just as for the chiral model of TBG, a flat band at zero for a given $\alpha$ means that
$\alpha\in\mathcal{A}\ \Longleftrightarrow\ \ \forall\,k\in\mathbb{C}\ \ker_{H^{2}(\mathbb{C}/\Lambda;\mathbb{C})}P(\alpha,k)\neq\{0\}.$ (2.6)
We denote the set of $\alpha$’s for which (2.6) holds by $\mathcal{A}$, the same notation used in the chiral model [Be*22]. It follows from [GaZw24] that $\mathcal{A}$ is a discrete subset of $\mathbb{C}$, and that
$\mathcal{A}=-\mathcal{A}=\overline{\mathcal{A}}.$
It is shown, together with the magic $\alpha$’s for the chiral model (2.1), both with $U$ given by (2.3), in Figure 4
Figure 4. The sets $\mathcal{A}$}, of magic $\alpha$’s for the chiral model (2.1) (in red) and the scalar model (2.4) (in blue) with $U$ given by (2.3). Higher multiplicities are indicated by additional circles: in the scalar model the real $\alpha$’s have multiplicity 22. When we interpolate between the chiral model and the scalar model, the multiplicity two real $\alpha$’s split and travel in opposite directions to become magic $\alpha$’s for the chiral model: see https://math.berkeley.edu/~zworski/Spec.mp4, where the magic $\alpha$’s are plotted as a function of $t\in[0,1]$ which linearly interpolates between the chiral and scalar models.
We define multiplicity as follows: if for $\alpha\in\Omega$, there is $k_{0}\in\mathbb{C}$ such that $\ker P(\alpha,k_{0})=\{0\}$, then we define
$m(\alpha,k):=\frac{1}{2\pi i}\operatorname{tr}\oint_{\partial D}P(\alpha,\zeta)^{-1}\partial_{\zeta}P(\alpha,\zeta)d\zeta,$ (2.7)
where the integral is over the positively oriented boundary of a disc $D$ which contains $k$ as the only possible pole of $\zeta\mapsto P(\alpha,\zeta)$. Otherwise, we put $m(\alpha,k)=\infty$ for all $k\in\mathbb{C}$.
2.2. Symmetries
To simplify notation, we write $P(k):=P(\alpha,k)$, where $P(\alpha,k)$ was defined in (1.1) with $V$ satisfying (2.5).
We define the following symmetries:
$\begin{gathered}\Omega u(z):=u(\omega z),\ \ \ \mathscr{L}_{\gamma}u(z)=u(z+\gamma),\ \ \ \mathscr{E}_{1}u(z)=u(-z),\\ \ \mathscr{A}_{1}u(z)=\overline{u(\bar{z})},\ \ \ \ \mathscr{R}_{1}u(z)=u(\bar{z}),\end{gathered}$
with the following commutation properties with $P(k)$: (here $\gamma\in\Lambda$ defined in (2.2))
$\begin{gathered}\mathscr{E}_{1}P(k)=P(-k)\mathscr{E}_{1},\ \ \ \mathscr{L}_{\gamma}P(k)=P(k)\mathscr{L}_{\gamma},\ \ \ P(k)\Omega=\omega\Omega P(\omega k),\\ \mathscr{A}_{1}P(k)=P(-\bar{k})\mathscr{A}_{1},\ \ \ \mathscr{R}_{1}P(\alpha,k)=P(\bar{\alpha},\bar{k})^{*}\mathscr{R}_{1}.\end{gathered}$ (2.8)
It is useful to combine $\mathscr{E}_{1}$ and $\Omega$:
$\mathscr{C}_{1}:=\mathscr{E}_{1}\Omega=\Omega\mathscr{E}_{1},\ \ \ P(k)\mathscr{C}_{1}=-\tau\mathscr{C}_{1}P(\tau k),\ \ \ \mathscr{C}_{1}u(z)=u(\tau z),\ \ \ \tau:=e^{-\pi i/3}.$ (2.9)
(Nothing is lost here since $\mathscr{C}^{3}_{1}=\mathscr{E}_{1}$ and $\mathscr{E}_{1}\mathscr{C}_{1}=\Omega$.)
We also note that for $\mathbf{a}\in\Lambda^{*}$
$P(k+\mathbf{a})=e^{-i\langle z,\mathbf{a}\rangle}P(k)e^{i\langle z,\mathbf{a}\rangle},\ \ u(z)\mapsto e^{i\langle z,\mathbf{a}\rangle}u(z)\text{ is unitary on $L^{2}(\mathbb{C}/\Lambda)$.}$
Using these symmetries, definition of the multiplicity (2.7) and [GaZw24, Theorem 1] we obtain
$m(\alpha,k)=\left\{\begin{array}[]{ll}\ \ \infty&\alpha\in\mathcal{A},\\ 2\operatorname{1\thinspace l}_{\Lambda^{*}}(k)&\alpha\notin\mathcal{A}.\end{array}\right$. (2.10)
In particular,
$\forall\alpha\in\mathbb{C},\ \ker P(\alpha,0)\neq\{0\}.$
We now relate $P(k)$ to the operator (1.2):
$D_{\rm{S}}=D_{\rm{S}}(\alpha)=\begin{pmatrix}2D_{\bar{z}}&\alpha V(z)\\ \alpha&2D_{\bar{z}}\end{pmatrix}:H^{1}(\mathbb{C}/\Lambda;\mathbb{C}^{2})\to L^{2}(\mathbb{C}/\Lambda;\mathbb{C}^{2}).$ (2.11)
We have
$(D_{\rm{S}}+k)\begin{pmatrix}u_{1}\\ u_{2}\end{pmatrix}=\begin{pmatrix}f\\ g\end{pmatrix}\ \Longleftrightarrow\ \left\{\begin{array}[]{l}\alpha u_{1}=(2D_{\bar{z}}+k)u_{2}-g,\\ P(k)u_{2}=\alpha f+(2D_{\bar{z}}+k)g.\end{array}\right$. (2.12)
Hence, with $\pi_{1}((u,v)^{t})=u$, $\pi_{2}((u,v)^{t})=v$, $\iota f:=(f,0)^{t}$,
[…]
For $\alpha\notin\mathcal{A}$, (2.13) gives (with integrals is over a positively oriented circle around 0)
$\begin{gathered}\Pi(\alpha):=\frac{1}{2\pi i}\oint_{0}(D_{S}(\alpha)+\zeta)^{-1}d\zeta,=\tfrac{1}{2}\begin{pmatrix}a(\alpha)&b(\alpha)\\ c(\alpha)&d(\alpha)\end{pmatrix},\\ a(\alpha):=\frac{1}{2\pi i}\oint\partial_{k}P(k)P(k)^{-1}dk,\ \ d(\alpha):=\frac{1}{2\pi i}\oint P(k)^{-1}\partial_{k}P(k)dk.\end{gathered}$ (2.14)
Hence
$\operatorname{tr}\Pi(\alpha)=\tfrac{1}{2}(\operatorname{tr}a(\alpha)+\operatorname{tr}d(\alpha))=\operatorname{tr}a(\alpha)=m(\alpha,0).$ (2.15)
(The equality of the traces of $a(\alpha)$ and $d(\alpha)$ follows from the Gohberg–Sigal theory – see for instance [DyZw19, Theorem C.11].) From (2.7) we conclude that
$(D_{S}(\alpha)+k)^{-1}=\frac{\Pi(\alpha)}{k}-\frac{D_{S}\Pi(\alpha)}{k^{2}}+A(k),$ (2.16)
and $k\mapsto A(k)$ is holomorphic near $k=0$.
We record the following symmetries of $D_{S}$: the translation symmetry $\mathscr{L}_{\gamma}D_{\rm{S}}=D_{\rm{S}}\mathscr{L}_{\gamma}$, and
$\mathscr{C}_{2}\begin{pmatrix}u_{1}\\ u_{2}\end{pmatrix}:=\begin{pmatrix}\bar{\tau}u_{1}(\tau\bullet)\\ u_{2}(\tau\bullet)\end{pmatrix},\ \ \ \mathscr{C}_{2}(D_{\rm{S}}(\alpha)+k)=\tau(D_{S}(\alpha)+\bar{\tau}k)\mathscr{C}_{2},\ \ \ \mathscr{C}_{2}\Pi=\Pi\mathscr{C}_{2},$ (2.17)
$\mathscr{R}_{2}\begin{pmatrix}u_{1}\\ u_{2}\end{pmatrix}:=\begin{pmatrix}u_{2}(\bar{\bullet})\\ u_{1}(\bar{\bullet})\end{pmatrix},\ \ \ \mathscr{R}_{2}(D_{\rm{S}}(\alpha)+k)=(D_{\rm{S}}(\bar{\alpha})+\bar{k})^{*}\mathscr{R}_{2},$
$\mathscr{A}_{2}\begin{pmatrix}u_{1}\\ u_{2}\end{pmatrix}:=\begin{pmatrix}-\bar{u}_{2}\\ \ \ \bar{u}_{1}\end{pmatrix},\ \ \ \mathscr{A}_{2}(D_{\rm{S}}(\alpha)+k)=-(D_{\rm{S}}(\alpha)-k)^{*}\mathscr{A}_{2}$
We note that
$\mathscr{R}_{2}\mathscr{C}_{2}=\bar{\tau}\mathscr{C}_{2}^{-1}\mathscr{R}_{2},\ \ \ \ \mathscr{A}_{2}\mathscr{C}_{2}=\tau\mathscr{C}_{2}\mathscr{A}_{2}.$ (2.18)
A self-adjoint operator associated to $D_{\rm{S}}$ is given by
$H_{\rm S}(\alpha,k):=\begin{pmatrix}0&D_{\rm{S}}(\alpha)+k\\ (D_{\rm{S}}(\alpha)+k)^{*}&0\end{pmatrix}:H^{1}(\mathbb{C}/\Lambda;\mathbb{C}^{4})\to L^{2}(\mathbb{C}/\Lambda;\mathbb{C}^{4})$
It enjoys the standard chiral symmetry
$\mathscr{W}H(\alpha,k)\mathscr{W}=-H(\alpha,k),\ \ \ \mathscr{W}\begin{pmatrix}\mathbf{u}\\ \mathbf{v}\end{pmatrix}:=\begin{pmatrix}-\mathbf{u}\\ \ \ \mathbf{v}\end{pmatrix},\ \ \ \mathbf{u},\mathbf{v}\in L^{2}(\mathbb{C}/\Lambda;\mathbb{C}^{2}).$ (2.19)
The symmetries of $P(k)$ give us also
$\mathscr{C}_{4}H(\alpha,k)=H(\alpha,\bar{\tau}k)\mathscr{C}_{4},\ \ \ \mathscr{C}_{4}\begin{pmatrix}\mathbf{u}\\ \mathbf{v}\end{pmatrix}:=\begin{pmatrix}\ \mathscr{C}_{2}\mathbf{u}\\ \tau\mathscr{C}_{2}\mathbf{v}\end{pmatrix},\ \ \ \mathscr{C}_{4}\mathscr{W}=\mathscr{W}\mathscr{C}_{4}.$ (2.20)
Finally,
$\mathscr{R}_{4}H(\alpha,k)=H(\bar{\alpha},\bar{k})\mathscr{R}_{4},\ \ \ \mathscr{R}_{4}\begin{pmatrix}\mathbf{u}\\ \mathbf{v}\end{pmatrix}:=\begin{pmatrix}\mathscr{R}_{2}\mathbf{v}\\ \mathscr{R}_{2}\mathbf{u}\end{pmatrix},\ \ \ \mathscr{C}_{4}\mathscr{R}_{4}=\mathscr{R}_{4}\mathscr{C}_{4}^{-1},$ (2.21)
$\mathscr{A}_{4}H(\alpha,k)=-H(\alpha,-k)\mathscr{A}_{4},\ \ \ \mathscr{A}_{4}\begin{pmatrix}\mathbf{u}\\ \mathbf{v}\end{pmatrix}:=\begin{pmatrix}\mathscr{A}_{2}\mathbf{v}\\ \mathscr{A}_{2}\mathbf{u}\end{pmatrix},\ \ \ \mathscr{C}_{4}\mathscr{A}_{4}=\mathscr{A}_{4}\mathscr{C}_{2},$ (2.22)
We also define (dropping $\mathbb{C}$ when $n=1$)
$L^{2}_{\ell}(\mathbb{C}/\Lambda,\mathbb{C}^{n}):=\{u\in L^{2}(\mathbb{C}/\Lambda,\mathbb{C}^{n}):\mathscr{C}_{n}u=\tau^{\ell}u\},\ \ \ n=1,2,4,$
noting that
$\begin{gathered}L^{2}_{\ell}(\mathbb{C}/\Lambda;\mathbb{C}^{2})=L^{2}_{\ell+1}(\mathbb{C}/\Lambda)\oplus L^{2}_{\ell}(\mathbb{C}/\Lambda),\\ L^{2}_{\ell}(\mathbb{C}/\Lambda;\mathbb{C}^{4})=L^{2}_{\ell}(\mathbb{C}/\Lambda;\mathbb{C}^{2})\oplus L^{2}_{\ell-1}(\mathbb{C}/\Lambda;\mathbb{C}^{2}),\\ \mathscr{R}_{1}:L^{2}_{\ell}(\mathbb{C}/\Lambda)\to L^{2}_{-\ell}(\mathbb{C}/\Lambda),\ \ \mathscr{R}_{2}:L^{2}_{\ell}(\mathbb{C}/\Lambda;\mathbb{C}^{2})\to L^{2}_{-\ell-1}(\mathbb{C}/\Lambda;\mathbb{C}^{2}),\\ \mathscr{R}_{4},\mathscr{A}_{4}:L^{2}_{\ell}(\mathbb{C}/\Lambda;\mathbb{C}^{4})\to L^{2}_{-\ell}(\mathbb{C}/\Lambda;\mathbb{C}^{2}).\end{gathered}$ (2.23)
We then have
$\begin{gathered}D_{\rm{S}}(\alpha):(H^{1}\cap L^{2}_{\ell})(\mathbb{C}/\Lambda;\mathbb{C}^{2})\to L^{2}_{\ell+1}(\mathbb{C}/\Lambda;\mathbb{C}^{2}),\\ H_{\ell}(\alpha):=H(\alpha):=H(\alpha,0):(H^{1}\cap L^{2}_{\ell})(\mathbb{C}/\Lambda;\mathbb{C}^{4})\to L^{2}_{\ell}(\mathbb{C}/\Lambda;\mathbb{C}^{4}).\end{gathered}$ (2.24)
3. Protected states and magic alphas in the scalar model
This section is devoted to the proof of Theorem 1. We first investigate the family of protected states, that is nontrivial elements of the kernel of $D_{\rm{S}}(\alpha)$ which exist for all $\alpha$. In the process we will define the discrete sets $\mathcal{A}$ and $\mathcal{B}$ which appear in Theorem 1. We then discuss the presence of Dirac points for $\alpha\in\mathcal{B}$ and the relation between
3.1. Family of protected states
We have the following result proved using methods similar to those in [Be*22], [BHZ24] and [Be*23] but with some new twists (no pun intended).
Proposition 1.
Let $\mathcal{A}$ be the set of magic $\alpha$’s defined in (2.6). There exists a discrete set $\mathcal{B}$}, $\mathcal{B}\cap\mathcal{A}=\emptyset$, such that for $\alpha\neq 0$,
$\dim\ker P(\alpha,0)=\dim\ker D_{\rm{S}}(\alpha)=\left\{\begin{array}[]{ll}1,&\alpha\notin\mathcal{A}\cup{\mathcal{B}},\\ 2,&\alpha\in{\mathcal{B}}.\end{array}\right$. (3.1)
Moreover, there exists a holomorphic family
$\mathbb{C}\ni\alpha\mapsto u(\alpha)\in\ker P(\alpha,0)\setminus\{0\},\ \ \ u(\alpha)\in L^{2}_{0}(\mathbb{C}/\Lambda;\mathbb{C}).$ (3.2)
Remarks. 1.The proposition and (2.13) show that for $\Pi(\alpha)$ defined in (2.14) we have
$D_{\rm{S}}(\alpha)\Pi(\alpha)\neq 0,\ \ \ \alpha\notin\mathcal{A}\cup{\mathcal{B}},$ (3.3)
that is, $D_{\rm{S}}(\alpha)$, has a Jordan block block structure at its eigenvalues (which are given by $\Lambda^{*}$. The way that $D_{\rm{S}}(\alpha)$ was defined we do not have (3.1) at $\alpha=0$. Instead, $\dim\ker P(0,0)=1$ and $\dim\ker D_{\rm{S}}(0)=2$, and from the point of view of Theorem 2, we should include 0 in $\mathcal{B}$. To have an agreement between the system and the scalar operator we could have considered
$\widetilde{D}_{S}(\alpha):=\begin{pmatrix}2D_{\bar{z}}&\alpha^{2}V(z)\\ 1&2D_{\bar{z}}\end{pmatrix}=\begin{pmatrix}1&0\\ 0&\alpha\end{pmatrix}^{-1}D_{S}(\alpha)\begin{pmatrix}1&0\\ 0&\alpha\end{pmatrix}$, (3.4)
To maintain a closer analogy to the chiral model (2.1) we keep $\alpha$ as a linear coupling constant.
2. As we will see, the set $\mathcal{B}$} has an interesting interpretation in terms of band structure: for $\alpha\in\mathcal{A}$, the operator $H(\alpha)$ has a flat band. For $\alpha\notin\mathcal{A}\cup{\mathcal{B}}$, the bands touch tangentially at 0, while for $\alpha\in{\mathcal{B}}$, the bands exhibit a Dirac point at 0. This will be explained in §3.2.
3. For $\alpha\notin\mathcal{A}\cup\mathcal{B}$, the unique solution to $P(\alpha,0)u=0$ satisfies symmetries inherited from the equation:
$u(z)=u(-z),\ \ \ u(\omega z)=u(z),\ \ \ \overline{u(\bar{z})}=u(z),\ \ \ u(z+\gamma)=u(z),\ \ \ \gamma\in\Lambda.$ (3.5)
We note that this also implies that (since $e^{\frac{1}{3}\pi i}=-\omega^{2}$)
$u(e^{\frac{1}{3}\pi i}z)=u(z),\ \ \ \ u(-\bar{z})=\overline{u(z)}.$
Analytic continuation shows that these properties remain true for the family $\mathbb{C}\ni\alpha\mapsto u(\alpha)$.
Proof.
We consider a family based on $\widetilde{D}_{S}(\alpha)$ defined in (3.4) and the corresponding self-adjoint family
$\widetilde{H}(\alpha):=\begin{pmatrix}0&\widetilde{D}_{\rm{S}}(\alpha)\\ \widetilde{D}_{\rm{S}}(\alpha)^{*}&0\end{pmatrix}$.
Away from 0 these are equivalent to $D_{S}(\alpha)$ and $H(\alpha)$ using (3.4). Since $\mathscr{C}_{2}$ commutes with the conjugation in (3.4) we still have the mapping properties (2.24).
We have
$\widetilde{H}(0)=\begin{pmatrix}&&2D_{\bar{z}}&0\\ &&1&2D_{\bar{z}}\\ 2D_{z}&1&&\\ 0&2D_{z}\end{pmatrix},\ \ \ \ker\widetilde{H}(0)=\{\mathbf{e}_{1},\mathbf{e}_{4}\},\ \ \ \mathbf{e}_{1}\in L^{2}_{-1},\ \ \mathbf{e}_{4}\in L^{2}_{1}.$
so that in the notation of (2.24)
$\dim\ker\widetilde{H}_{\ell}(0)=\left\{\begin{array}[]{ll}1&\ell=\pm 1\\ 0&\text{otherwise. }\end{array}\right$. (3.6)
In view of (2.20), $\mathcal{W}\widetilde{H}_{\ell}(\alpha)\mathcal{W}=-\widetilde{H}_{\ell}(\alpha)$, and hence the spectrum of $\widetilde{H}_{\ell}(\alpha)$ is symmetric about 0. At $\alpha=0$, $0\in\operatorname{Spec}(\widetilde{H}_{\ell}(0))$ only for $\ell=\pm 1$, and then it is a simple eigenvalue. Hence, the multiplicity of 0 as an eigenvalue of $\widetilde{H}_{\ell}(\alpha)$, $\ell=\pm 1$ has to be odd, while that $\widetilde{H}_{\ell}(\alpha)$, $\ell\notin\{\pm 1\}$ has to be even. In particular,
$\dim\ker\widetilde{H}_{\pm 1}(\alpha)\geq 1.$
In view of (2.24)
$\begin{gathered}\ker\widetilde{H}_{\ell}(\alpha)=\ker\widetilde{D}_{\rm{S}}(\alpha)^{*}|_{L^{2}_{\ell}}\oplus\{0\}+\{0\}\oplus\ker\widetilde{D}_{\rm{S}}(\alpha)|_{L^{2}_{\ell-1}},\\ \widetilde{D}_{\rm{S}}(\alpha)|_{L^{2}_{\ell-1}}:L^{2}_{\ell-1}\to L^{2}_{\ell},\ \ \ \widetilde{D}_{\rm{S}}(\alpha)^{*}|_{L^{2}_{\ell}}=(\widetilde{D}_{\rm{S}}(\alpha)|_{L^{2}_{\ell-1}})^{*}=L^{2}_{\ell}\to L^{2}_{\ell-1}.\end{gathered}$ (3.7)
Since it is true at $\alpha=0$, the continuity of (simple) eigenfunctions shows that for $|\alpha|\ll 1$,
$\ker\widetilde{H}_{1}(\alpha)=\{0\}\oplus\ker\widetilde{D}_{\rm{S}}(\alpha)|_{L^{2}_{0}},$
and
$\ker\widetilde{H}_{-1}(\alpha)=\mathscr{A}_{4}\ker\widetilde{H}_{1}(\alpha)=\ker\widetilde{D}_{\rm{S}}(\alpha)^{*}|_{L^{2}_{-1}}\oplus\{0\}.$
This also shows that
$\dim\ker\widetilde{D}_{S}(\alpha)=1,\ \ \ |\alpha|\ll 1,$
as $\dim\ker\widetilde{D}_{S}(\alpha)$ is upper semicontinuous (see for instance [Hö88, Theorem 2.5.3]). Since ${\rm{rank}}\,\Pi(\alpha)=2$ (this follows from (2.15)), we have $\widetilde{D}_{\rm{S}}(\alpha)\Pi(\alpha)\neq 0$ for $|\alpha|\ll 1$, that is,
$\ker\widetilde{D}_{\rm{S}}(\alpha)=\widetilde{D}_{\rm{S}}(\alpha)\Pi(\alpha)L^{2}.$
In view of the mapping properties in (2.24), we conclude that $\Pi(\alpha)=\mathbb{C}u(\alpha)\oplus\mathbb{C}v(\alpha)$, where
$u(\alpha)\in\ker\widetilde{D}_{\rm{S}}(\alpha)\subset L^{2}_{0}(\mathbb{C}/\Lambda;\mathbb{C}^{2}),\ \ v(\alpha)\in L^{2}_{-1}(\mathbb{C}/\Lambda;\mathbb{C}^{2}),\ \ |\alpha|\ll 1.$
In particular, if $\pi_{\ell}:L^{2}\to L^{2}_{\ell}$ are the orthogonal projections, we have for $|\alpha|\ll 1$,
$(\pi_{0}+\pi_{-1})\Pi(\alpha)=\Pi(\alpha).$ (3.8)
We now note that for $\mathbb{C}\setminus\mathcal{A}\ni\alpha\mapsto\Pi(\alpha)$ is a holomorphic family of projections (this follows from applying $\partial_{\bar{\alpha}}$ to the definition of $\Pi(\alpha)$ in (2.14)). We then see that (3.8) holds for $\alpha\in\mathbb{C}\setminus\mathcal{A}$.
Since $\mathbb{C}\setminus\mathcal{A}\ni\alpha\mapsto\operatorname{tr}\pi_{0}\Pi(\alpha)\pi_{0}$ is holomorphic and hence identically equal to $1$, and since $\pi_{0}\Pi(\alpha)\pi_{0}=\pi_{0}\Pi(\alpha)$ (the projections $\pi_{\ell}$ commute with $\Pi(\alpha)$ thanks to (2.17)), we conclude that
$\dim\pi_{0}\Pi(\alpha)L^{2}=1,\ \ \alpha\in\mathbb{C}\setminus\mathcal{A}.$ (3.9)
Hence, if (3.1) does not hold then
$\ker\widetilde{D}_{S}(\alpha)|_{L^{2}_{-1}}\neq\{0\},$
that is, there exists a non-zero $v\in L^{2}_{-1}(\mathbb{C}/\Lambda)$ such that $P(\alpha,0)v=0$ (we then have $2D_{\bar{z}}v,v]^{t}\in\ker\widetilde{D}_{\rm{S}}|_{L^{2}_{-1}(\mathbb{C}/\Lambda;\mathbb{C}^{2})}$). Since
$D_{\bar{z}}^{2}:L^{2}_{-1}(\mathbb{C}/\Lambda)\to L^{2}_{1}(\mathbb{C}/\Lambda)$
is invertible, and $u\mapsto Vu$, takes $L^{2}_{-1}\to L^{2}_{1}$, we have
$P(\alpha,0)|_{L^{2}_{-1}}=(2D_{\bar{z}})^{2}|_{L^{2}_{-1}}(I-\alpha^{2}K),\ \ \ K:=[(2D_{\bar{z}})^{2}|_{L^{2}_{-1}}]^{-1}V:L^{2}_{-1}\to L^{2}_{-1}.$
The operator $K$ is compact and hence $I-\alpha^{2}K$ is invertible outside of a discrete set of $\alpha$’s. We define the set $\mathcal{B}$ as follows
$\alpha\in\mathcal{B}\ \Longleftrightarrow\ \ker P(\alpha,0)|_{L^{2}_{-1}}\neq\{0\}\ \text{ and }\alpha\notin\mathcal{A}.$ (3.10)
We note that for $\alpha\in\mathcal{B}$, $\dim\ker\widetilde{D}_{\rm{S}}(\alpha)=2$. In fact, if there existed three independent solutions $\widetilde{D}_{\rm{S}}(\alpha)u_{j}=0$, for any $z_{0}\in\mathbb{C}/\Lambda$ we could find $\gamma\in\mathbb{C}^{3}\setminus\{0\}$ such that $\sum_{j=1}^{3}\gamma_{j}u_{j}(z_{0})=0$ (since $u_{j}(z_{0})\in\mathbb{C}^{2}$ these are two equations). Since $\sum_{j=1}^{3}\gamma_{j}u_{j}\neq 0$ (from the independence of $u_{j}$’s) the now standard argument (see [TKV19] and [Zw24, §6]) shows that $\alpha\in\mathcal{A}$.
To see (3.2) we modify the proof of [BHZ24, Proposition 2.3]. For that we fix $\alpha$ and define a new operator
$\widetilde{H}(\zeta):=\begin{pmatrix}0&\widetilde{D}_{\rm{S}}(\alpha+\zeta)\\ \widetilde{D}_{\rm{S}}(\alpha+\bar{\zeta})^{*}&0\end{pmatrix}$,
which has the property that $\widetilde{H}(\zeta)=\widetilde{H}(\bar{\zeta})^{*}$. We can restrict this operator to $L^{2}_{1}(\mathbb{C}/\Lambda;\mathbb{C}^{4})$ and, as we already know that $\ker\widetilde{D}_{S}(\alpha+\zeta)|_{L^{2}_{0}}$ is nontrivial, $\dim\ker\widetilde{H}_{1}(\zeta)\geq 1$. Rellich’s theorem [Ka80, Chapter VII, Theorem 3.9], then shows that there exists a holomorphic family ${\rm neigh}_{\mathbb{C}}(0)\ni\zeta\to\widetilde{u}(\zeta)\in\ker\widetilde{H}_{1}(\zeta)\setminus\{0\}$ (it corresponds to the 0 eigenfunctions of $\widetilde{H}_{1}(\zeta)$). If $\alpha+\zeta\notin\mathcal{A}$, then $\widetilde{u}(\zeta)=(0,\widetilde{\mathbf{u}}(\zeta))$, $\ker\widetilde{D}_{\rm{S}}(\alpha+\zeta)|_{L^{2}_{0}}=\mathbb{C}\widetilde{\mathbf{u}}(\zeta)$. This construction is local near any $\alpha\in\mathbb{C}$ and by partition of unity we can find a smooth family $\alpha\mapsto\mathbf{u}_{1}(\alpha)\in\ker\widetilde{D}_{\rm{S}}(\alpha)|_{L^{2}_{0}}$.
To modify to obtain a holomorphic family we first note that
$0=\partial_{\bar{\alpha}}(\widetilde{D}_{\rm{S}}(\alpha)\mathbf{u}_{1}(\alpha))=\widetilde{D}_{\rm{S}}(\alpha)(\partial_{\bar{\alpha}}\mathbf{u}_{1}(\alpha))\ \Longrightarrow\ \partial_{\bar{\alpha}}\mathbf{u}_{1}(\alpha)\in\ker\widetilde{D}_{\rm{S}}(\alpha)|_{L^{2}_{0}}.$
Using the fact that for $\alpha\notin\mathcal{A}$ this kernel is one dimensional we see that
$\partial_{\bar{\alpha}}\mathbf{u}_{1}(\alpha)=f(\alpha)\mathbf{u}_{1}(\alpha),\ \ \ f(\alpha):=\frac{\langle\partial_{\bar{\alpha}}\mathbf{u}_{1}(\alpha),\mathbf{u}_{1}(\alpha)\rangle}{\|\mathbf{u}_{1}(\alpha)\|^{2}},\ \ \ \alpha\notin\mathcal{A}.$ (3.11)
The formula for $f(\alpha)$ defines a smooth function $f\in C^{\infty}(\mathbb{C})$ and that means that the first formula in (3.11) holds everywhere. The equation $\partial_{\bar{\alpha}}F(\alpha)=f(\alpha)$ (see for instance [HöI, Theorem 4.4.6] applied with $P=\partial_{\bar{\alpha}}$ and $X=\mathbb{C}$) can be solved with $F\in C^{\infty}(\mathbb{C})$. This shows that $\mathbf{u}(\alpha):=\exp(-F(\alpha))\mathbf{u}_{1}(\alpha)$ is holomorphic. Since $\mathbf{u}(\alpha)=[2D_{\bar{z}}u(\alpha),u(\alpha)]$, we obtain (3.2). ∎
Remark. For $\alpha\in\mathbb{R}\setminus\{0\}$, having a kernel of $D_{\rm{S}}(\alpha)|_{L^{2}_{-1}}$ is a natural possibility for another reason than the one presented before (3.10). It implies that $\ker H_{0}(\alpha)\neq 0$. Since $\ker H_{\ell}(0)=0$, $\ell\neq\pm 1$, the chiral symmetry shows that $\dim\ker H_{\ell}(\alpha)\equiv 0\mod 2$ and if $\dim\ker H(\alpha)>2$, then there exists a unique (in view of (3.9)) $\ell\neq\pm 1$ such that $\dim\ker H_{\ell}(\alpha)=2$ (For $\alpha\in\mathbb{R}\setminus\mathcal{A}$, we have $\dim\ker H(\alpha)\leq 4$, as otherwise $\mathscr{R}_{4}H(\alpha)=H(\alpha)\mathscr{R}_{4}$ would show that $\dim\ker D_{\rm{S}}(\alpha)>2$ contradicting (2.15).) Since $\mathscr{R}_{4}:\ker H_{\ell}(\alpha)\to H_{-\ell}(\alpha)$, that means that $\ell\equiv-\ell\mod 6$, that is $\ell=0$ or $\ell=3$. The case $\ell=0$ corresponds to a nontrivial kernel of $D_{\rm{S}}(\alpha)|_{L^{2}_{-1}}$ (see (2.24)). The case $\ell=3$ is excluded by the assumption that $\alpha\notin\mathcal{A}$. In fact, we would then have a non-trivial kernel of $D_{\rm{S}}(\alpha)|_{L^{2}_{2}}$. But that would mean that we would have $v$, with $D_{\rm{S}}(\alpha)v=0$, $\bar{\tau}v_{1}(\tau z),v_{2}(\tau z)]^{t}=[\tau^{2}v_{1}(z),\tau^{2}v_{2}(z)]^{t}$, that is $v_{1}(\tau z)=-v_{1}(z)$ and $v_{2}(\tau z)=\tau^{2}v_{2}(z)$. In particular $v(z)=0$ and the theta function argument (see [TKV19] and [Zw24, §6]) shows that $\alpha\in\mathcal{A}$ (see Proposition 2).
3.2. Dirac points for a discrete set of $\alpha$’s
Suppose that $\alpha\in{\mathcal{B}}$ defined in (3.10). Then $0\in\operatorname{Spec}D_{\rm{S}}(\alpha)$ has geometric (and algebraic) multiplicity $2$ and there exist
$\mathbf{u}\in\ker D_{\rm{S}}(\alpha)|_{L^{2}_{0}},\ \ \ \mathbf{v}\in\ker D_{\rm{S}}(\alpha)|_{L^{2}_{-1}},\ \ \langle\mathbf{u},\mathbf{v}\rangle=0,\ \ \ \|\mathbf{u}\|=\|\mathbf{v}\|=1.$ (3.12)
We consider the Wronskian of $\mathbf{u}$ and $\mathbf{v}$,
$\operatorname{Wr}(\mathbf{u},\mathbf{v})=u_{1}v_{2}-u_{2}v_{1},$ (3.13)
and note that it is non-zero as otherwise either $\mathbf{u}$ or $\mathbf{v}$ would have a zero at 0 and then $\alpha\in\mathcal{A}$ – see Proposition 2. We can then set up a Grushin problem for $H(\alpha,0)$ which remains invertible when we replace $H(\alpha,0)$ by $H(\alpha,k)$ for small $k$:
[…]
Then
$\mathscr{H}(\alpha,k)^{-1}=\begin{pmatrix}E(\alpha,k)&E_{+}(\alpha,k)\\ E_{-}(\alpha,k)&E_{-+}(\alpha,k)\end{pmatrix},\ \ E_{\pm}(\alpha,0)=R_{\pm},$
where (see [TaZw23, Proposition 2.12])
$\begin{split}E_{-+}(\alpha,k)&=-R_{+}\begin{pmatrix}0&k\\ \bar{k}&0\end{pmatrix}R_{-}+\mathcal{O}(|k|^{2})\\ &=\begin{pmatrix}0&\bar{k}A^{*}\\ kA&0\end{pmatrix}+\mathcal{O}(|k|^{2}),\ \ \ A:=\tfrac{\sqrt{3}}{2}W\begin{pmatrix}\ \ 0&1\\ -1&0\end{pmatrix}\end{split}$ (3.15)
where $W=W(\mathbf{u},\mathbf{v})$. This follows from the following calculations:
$\begin{split}&\langle\mathbf{u},\mathscr{A}_{2}\mathbf{u}\rangle=\int_{\mathbb{C}/\Lambda}u_{1}\overline{(-\bar{u}_{2})}+u_{2}\overline{(\bar{u}_{1})}dm(z)=0,\ \ \ \langle\mathbf{v},\mathscr{A}_{2}\mathbf{v}\rangle=0,\\ \ \ \ &\langle\mathbf{u},\mathscr{A}_{2}\mathbf{v}\rangle=\int_{\mathbb{C}/\Lambda}u_{1}\overline{(-\bar{v}_{2})}+u_{2}\overline{(\bar{v}_{1})}dm(z)=\tfrac{\sqrt{3}}{2}\operatorname{Wr}(\mathbf{v},\mathbf{u})=-\langle\mathbf{v},\mathscr{A}_{2}\mathbf{u}\rangle,\end{split}$
and
$\begin{split}\begin{pmatrix}0&k\\ \bar{k}&0\end{pmatrix}R_{-}u_{-}&=\begin{pmatrix}0&k\\ \bar{k}&0\end{pmatrix}\left(u_{-,1}\begin{pmatrix}0\\ \mathbf{u}\end{pmatrix}+u_{-,2}\begin{pmatrix}0\\ \mathbf{v}\end{pmatrix}+u_{-,3}\begin{pmatrix}\mathscr{A}_{2}\mathbf{u}\\ 0\end{pmatrix}+u_{-,3}\begin{pmatrix}\mathscr{A}_{2}\mathbf{v}\\ 0\end{pmatrix}\right)\\ &=u_{-,1}\begin{pmatrix}k\mathbf{u}\\ 0\end{pmatrix}+u_{-,2}\begin{pmatrix}k\mathbf{v}\\ 0\end{pmatrix}+u_{-,3}\begin{pmatrix}0\\ \bar{k}\mathscr{A}_{2}\mathbf{u}\end{pmatrix}+u_{-,3}\begin{pmatrix}0\\ \bar{k}\mathscr{A}_{2}\mathbf{v}\end{pmatrix}\end{split}$
To study the eigenvalues near 0 of $H(\alpha,k)$ for $|k|\ll 1$ we observe that the Grushin problem (3.14) is well posed when $H(\alpha,k)$ is replaced by $H(\alpha,k)-\lambda$, $|\lambda|\ll 1$. Using [TaZw23, Proposition 2.12] we then see that small eigenvalues of $H(\alpha,k)$ are given by the zeros of $\lambda\to\det(E^{\lambda}_{-+}(\alpha,k))$ where, in the notation of (3.15),
$\begin{split}E^{\lambda}_{-+}(\alpha,k)&=E_{-+}(\alpha,k)-\lambda E_{-}(\alpha,k)E_{+}(\alpha,k)+\mathcal{O}(\lambda^{2})\\ &=\begin{pmatrix}0&\bar{k}A^{*}\\ kA&0\end{pmatrix}-\lambda+\mathcal{O}(\lambda^{2}+|k|^{2}).\end{split}$
The eigenvalues of the first term on the right hand side are double and given by $\pm\frac{\sqrt{3}}{2}|k||W|$ and hence (with the labelling convention of (1.5))
$E_{\pm j}(\alpha,k)=\pm\tfrac{\sqrt{3}}{2}|\operatorname{Wr}(\mathbf{u},\mathbf{v})||k|+\mathcal{O}(|k|^{2}),\ \ \ \alpha\in{\mathcal{B}},\ \ j=1,2,$ (3.16)
while all the other eigenvalues are non-zero. As in [TKV19], the Wronskian is interpreted as the Fermi velocity, that is the slope of the Dirac cone. However, it occurs only for a discrete set of $\alpha$ at which the cone is double.
When $\alpha\notin{\mathcal{B}}\cup\mathcal{A}$, then $\ker D_{S}(\alpha)=\mathbb{C}\mathbf{u}$, and setting up the Grushin problem using $(0,\mathbf{u})^{t}$ and $(\mathscr{A}_{2}\mathbf{u},0)^{t}$, shows that
$E_{\pm 1}(\alpha,k)=\mathcal{O}(|k|^{2}),\ \ |E_{\pm 2}(\alpha,k)|>0,\ \ \ \alpha\notin{\mathcal{B}}\cup\mathcal{A}.$ (3.17)
3.3. Zeros of $z\mapsto u(\alpha,z)$
We start with some general considerations and first prove the following fact.
Proposition 2.
In the notation of (3.2) let $\mathbf{u}(\alpha)=[2D_{\bar{z}}u(\alpha),u(\alpha)]^{t}$. Then
$\alpha\in\mathcal{A}\ \Longleftrightarrow\ \exists\,z_{0}\in\mathbb{C}/\Lambda\ \mathbf{u}(z_{0})=0.$ (3.18)
Proof.
[…]
$\operatorname{Wr}(z,\bar{z})=e^{\frac{i}{2}(k\bar{z}+\bar{k}z)}w(z),\ \ \ w\in\mathscr{O}(\mathbf{C}).$
Since $\operatorname{Wr}$ is periodic and $z\mapsto e^{\frac{i}{2}(k\bar{z}+\bar{k}z)}$ is a bounded function, $w$ has to be constant, and for $k\notin\Lambda^{*}$ that constant has to vanish.
Since $\mathbf{u}$ is not identically zero and real analytic (it solves an elliptic equation with analytic coefficients – see [HöI, Theorem 8.6.1]), $\Omega:=\complement\mathbf{u}^{-1}(0)$ is an open dense set. The vanishing of $\operatorname{Wr}$ means that $\mathbf{u}$ and $\mathbf{v}$ are parallel and hence there exists $F\in C^{\infty}(\Omega)$ such that $\mathbf{v}=F\mathbf{u}$. Applying $2D_{\bar{z}}$ to both sides shows that $2D_{\bar{z}}F=-kF$. Hence for some $f\in C^{\infty}(\Omega)$,
$\mathbf{v}(z,\bar{z})=e^{-\frac{i}{2}(k\bar{z}+\bar{k}z)}f(z)\mathbf{u}(z,\bar{z}),\ \ \ \ \partial_{\bar{z}}f|_{\Omega}=0.$ (3.19)
As in the proof of [BHZ23a, (4.4)] we see that $f$ is meromorphic. In fact, for a fixed $z_{1}$ we put
$G_{0}(\zeta,\bar{\zeta}):=v_{1}(z,\bar{z})|_{z=z_{1}+\zeta},\ \ \ G_{j}(\zeta,\bar{\zeta}):=e^{-\frac{i}{2}(k\bar{z}+\bar{k}z)}u_{1}(z,\bar{z})|_{z=z_{1}+\zeta}.$
Real analyticity of $\mathbf{u}$ and $\mathbf{v}$ shows that $G_{j}\in\mathscr{O}(B_{\mathbb{C}^{2}}(0,\delta))$. With $g(\zeta):=f(z_{1}+\zeta)$, we have
$G_{0}(\zeta,\xi)=g(\zeta)G_{1}(\zeta,\xi),\ \ \ z_{1}+\zeta\in\Omega,\ \$
Now choose $\xi_{0}\in B_{\mathbb{C}}(0,\delta/2)$ such that $\zeta\mapsto G_{1}(\zeta,\xi_{0})$ is not identically zero (if no such $\xi_{0}$ existed, $G_{1}\equiv 0$, and hence, from the equation, $\mathbf{u}\equiv 0$). But then $\zeta\mapsto g(\zeta):=G_{1}(\zeta,\xi_{0})/G_{2}(\zeta,\xi_{0})$ is meromorphic near $\zeta=0$ and, as $z_{1}$ was arbitrary $f$ is meromorphic everywhere.
For $\mathbf{v}$ to be periodic, $f$ cannot be constant and hence it has to have poles. But that means that $\mathbf{u}$ has to have zeros. ∎
We have the following refinement of Proposition 1. To formulate it, we define the notion of multiplicity:
Definition 3.
For a solution $\mathbf{u}\in\ker D_{\rm{S}}(\alpha)$ we define the multiplicity of a zero at $z_{0}$ as
$M_{\mathbf{u}}(z_{0}):=\max\{m:[\partial_{z}^{m-1}\mathbf{u}](z_{0})=0\},$ (3.20)
with the convention that $M_{\mathbf{u}}(z_{0})=0$ if $\mathbf{u}(z_{0})\neq 0$.
Since $\mathbf{u}=[2D_{\bar{z}}u,u]^{t}$, $P(\alpha)u=0$, and $m=M_{\mathbf{u}}(z_{0})$ is equivalent to $u(z)=(z-z_{0})^{m}v(z)$, $v(z_{0})\neq 0$, $v\in C^{\infty}({\rm neigh}_{\mathbb{C}}(z_{0}))$. We note here that $\mathbf{u}(z_{0})=0$ and $D_{\rm{S}}(\alpha)\mathbf{u}(z)=0$ imply that $\partial_{\bar{z}}^{k}\mathbf{u}(z_{0})=0$ for all $k$. That means $\partial_{\bar{z}}^{k}u(z_{0})=0$ for all $k$ and hence we can factor out a positive power of $(z-z_{0})$.
Proposition 4.
In the notation of (3.2) let $\mathbf{u}(\alpha)=[2D_{\bar{z}}u(\alpha),u(\alpha)]^{t}$. For $\alpha\in\mathcal{A}$
$\dim\ker D_{\rm{S}}(\alpha)=\sum_{z\in\mathbb{C}/\Lambda}M_{\mathbf{u}(\alpha)}(z)=:m(\alpha).$ (3.21)
This means that for all $k\in\mathbb{C}/\Lambda^{*}$,
$\dim\ker H(\alpha,k)=2m(\alpha),$
which is the number of flat bands at 0.
Proof.
We prove that for $\alpha\in\mathcal{A}$, $\dim\ker D_{\rm{S}}(\alpha)\geq m(\alpha)$. For that suppose first that $\mathbf{u}(\alpha)$ has distinct zeros $z_{0},\cdots,z_{m-1}$ each of multiplicity one. We can assume that $m\geq 2$ as otherwise there is nothing to prove. We then choose
$w_{j}\notin\{z_{j}\}_{j=0}^{m-1},\ \ \ \ \ \sum_{j=0}^{m-1}w_{j}\equiv\sum_{j=0}^{m-1}z_{j},$
and put
$\mathbf{u}_{j}(z,\bar{z})=\frac{\theta(z-w_{0})\theta(z-w_{j})}{\theta(z-z_{0})\theta(z-z_{j})}\mathbf{u}(z,\bar{z}).$
The factor in front of $\mathbf{u}$ is a meromorphic function with simple poles at $z_{0}$ and $z_{j}$ (see [BHZ24, §3.1] for the review of properties of $\theta$ in a related context) but the properties of the zeros show that $\mathbf{u}_{j}\in\ker D_{\rm{S}}(\alpha)$. The functions $\mathbf{u}_{j}$ are linearly independent as
$c_{0}+\sum_{j=1}^{n-1}c_{j}\frac{\theta(z-w_{0})\theta(z-w_{j})}{\theta(z-z_{0})\theta(z-z_{j})}\equiv 0\ \ \Longrightarrow\ c_{0}=c_{1}=\cdots=c_{m-1}=0.$
(Put $z=z_{j}$, $j>0$ to see that $c_{j}=0$ for $j>0$.) The case of higher multiplicities is dealt with similarly but using meromorphic functions with poles of higher order. This proves that $\dim\ker D_{S}(\alpha)\geq m(\alpha)$.
To prove the opposite inequality, suppose that $\mathbf{u}$, $\mathbf{v}_{j}$, $1\leq j\leq M-1$, span $\ker D_{\rm{S}}(\alpha)$, $M=\dim\ker D_{\rm{S}}(\alpha)$. Then the Wronskians (3.13), $\operatorname{Wr}(\mathbf{u},\mathbf{v}_{j})=0$ as $\mathbf{u}$ has to vanish somewhere by Proposition 2 and the Wronskian is constant. But as in (3.19),
$\mathbf{v}_{j}(z,\bar{z})=f_{j}(z)\mathbf{u}(z,\bar{z}),\ \text{ \ \ $f_{j}(z)$, $1\leq j\leq m-1$, meromorphic,}$ (3.22)
$f_{0}\equiv 1$ and $\{f_{j}\}_{j=0}^{M-1}$ linearly independent. The span of $f_{j}$’s is contained in the space of meromorphic functions [^1] For an accessible introduction see Terry Tao’s blog https://terrytao.wordpress.com/2018/03/28/246c-notes-1-meromorphic-functions-on-riemann-surfaces-and-the-riemann-roch-theorem/ and the version of the Riemann–Roch theorem needed here is in formula (7) there. $L(D)$ where $D$ is the divisor defined by the zeros of $\mathbf{u}$ included according to their multiplicities defined in (3.20). Since $\dim L(D)=\deg D=m(\alpha)$ (see (3.21)), this means that the number of zeros of $\mathbf{u}$ has to be greater than or equal to $M$, the dimension of $\ker D_{\rm{S}}(\alpha)$. ∎
Remarks. 1. The same argument applies in the chiral model and settles [Zw24, Problem 14] – see [BHZ23b, Theorem 1].
2. Although we invoked a basic version of the Riemann–Roch theorem the proof that the number of poles of $f_{j}$’s has to be greater than $M$ is explicit. If the poles are all simple then
$f_{j}(z)=\sum_{k=1}^{K(j)}\lambda_{k}(j)\frac{\theta^{\prime}(z-a_{k}(j))}{\theta(z-a_{k}(j))}+c(j),\ \ \ \ \sum_{k=1}^{K(j)}\lambda_{k}(j)=0.$
For $f_{0}\equiv 1$, $f_{1},\cdots,f_{M-1}$ to be linearly independent we need the number of distinct zeros $\{a_{k}(j):1\leq k\leq K(j),1\leq j\leq M-1\}$ to be $\geq M$. It is not difficult to modify this construction to the case of poles of higher order.
Figure 5. When multiplicity of the band is one, the only possible zero of $\mathbf{u}$ is at 0 and that is illustrated in this figure. For the potential (2.2) in the scalar model with $V(z)=U_{\rm{BM}}(z)U_{\rm{BM}}(-z)$ this occurs only for complex $\alpha$’s.
3.4. Multiplicity two
Suppose that $\dim\ker D_{S}(\alpha)=2$ and $\alpha\in\mathcal{A}$. We remark that numerical investigation in the case of the Bistritzer–MacDonald potential (2.3) (that is, V(z)=U(z)U(z)V(z)=U(z)U(-z) with that $U$), this is the case for $\mathcal{A}\cap\mathbb{R}$.
The assumption $\dim\ker D_{S}(\alpha)=2$ means that $\dim\ker H(\alpha,0)=4$ and consequently $\dim\ker H(\alpha,k)=4$ for all $k$. Since $\dim\ker H_{\ell}(0,0)\equiv\dim\ker H_{\ell}(\alpha,0)\!\!\mod 2$, in view of (3.6), there exists $\ell\neq\pm 1$ such that $\dim\ker H_{\ell}(\alpha,0)=2$. But then (2.22) and (2.24) imply that $\ell\equiv-\ell\!\!\mod 6$, that is, $\ell=0,3$.
The kernel of $D_{\rm{S}}(\alpha)$ is consequently spanned by $\mathbf{u}\in L^{2}_{0}(\mathbb{C}/\Lambda;\mathbb{C}^{2})$ and $\mathbf{v}\in L^{2}_{\ell}(\mathbb{C}/\Lambda;\mathbb{C}^{2})$, $\ell$ is either 1-1 (the case of $\dim\ker H_{0}(\alpha,0)=2$) or $2$ (the case of $\dim\ker H_{3}(\alpha,0)=2$). The Wronskian (3.13), $\operatorname{Wr}(\mathbf{u},\mathbf{v})=0$ as otherwise we would have (3.16) and no flat bands. We note that $u_{1}(\pm z_{S})=u_{1}(0)=0$, and $v_{2}(\pm z_{S})=v_{2}(0)=0$ (for both $\ell=-1$ and $\ell=2$). Hence
$u_{2}(\pm z_{S})v_{1}(\pm z_{S})=u_{2}(0)v_{1}(0)=0.$
Hence, we have two exclusive cases ($\mathbf{u}$ cannot vanish at three points as then Proposition 4 would contradict the multiplicity two assumptions):
$\mathbf{u}(0)=0\ \text{ or }\ \mathbf{u}(\pm z_{S})=0$ (3.23)
Suppose first that $\mathbf{u}(0)=0$ and $\mathbf{u}(\pm z_{S})\neq 0$. Then $\mathbf{v}(\pm z_{S})=0$, and we can use the following well known
Lemma 5.
The unique (up to a multiplicative constant) meromorphic function with two poles at $\pm z_{S}$ on $\mathbb{C}/\Lambda$ and a (double) zero at 0 satisfies
$F(\tau z)=\tau^{2}F(z),\ \ \ \tau=e^{-\pi i/3}.$ (3.24)
Proof.
We first note that $F$ is unique as for any periodic meromorphic function the sum of the poles is congruent to the sum of the zeros and the sum of residues adds up to 0. We can then given an explicit formula for $F$:
[…]
This shows absolute convergence and F(z)=F(z)F(z)=F(-z). Hence to obtain (3.24) it suffices to show that $F(\omega z)=\bar{\omega}F(z)$. In fact, $\omega z_{S}=z_{S}+\gamma_{0}$, $\gamma_{0}\in\Lambda$, and
$\begin{split}F(\omega z)&=\bar{\omega}\sum_{\gamma\in\Lambda}\frac{1}{z+\gamma-\gamma_{0}-z_{S}}-\frac{1}{z+\gamma+\gamma_{0}+z_{S}}-\frac{1}{\gamma-\gamma_{0}-z_{S}}+\frac{1}{\gamma+\gamma_{0}+z_{S}}\\ &=\bar{\omega}F(z)+\sum_{\gamma\in\Lambda}\frac{1}{z+\gamma+z_{S}}-\frac{1}{z+\gamma+2\gamma_{0}+z_{S}}-\frac{1}{\gamma+z_{S}}+\frac{1}{\gamma+2\gamma_{0}+z_{S}}.\end{split}$
The sum on the right hand side vanishes as it converges uniformly and defines a periodic holomorphic function vanishing at 0. This completes the proof. ∎
Remark. The function $F$ can be expressed using theta functions (see [BHZ24, §3.1] for the definitions in the this context)
$F(z)=\frac{\theta(z)^{2}}{\theta(z-z_{S})\theta(z+z_{S})},$
and the conclusion follows from [BHZ24, (3.1)] and the fact that $\bar{\omega}z_{S}=z_{S}+\gamma_{0}$, $\gamma_{0}\in\Lambda$.
Since $F(z)\mathbf{v}$ is not a scalar multiple of $\mathbf{v}$, we have $\mathbf{u}=c_{0}F(z)\mathbf{v}\in L^{2}_{\ell+2}$. But for $\ell=-1,2$, $\ell+2\not\equiv 0\mod 6$ which is a contradiction. Hence only the second option in (3.23) is allowed. But then (3.24) implies that $\mathbf{v}=c_{0}F(z)\mathbf{u}\in L^{2}_{2}$. We summarise this with
Proposition 6.
The multiplicity of $\alpha\in\mathcal{A}$ is 2 (that is, $\dim\ker H(\alpha,k)=4$ for all $k$) if and only if $\mathbf{u}(\alpha)=[2D_{\bar{z}}u(\alpha),u(\alpha)]^{t}$ in (3.2) vanishes at $\pm z_{S}$. In that case,
$\ker D_{S}(\alpha)=\mathbb{C}\mathbf{u}\oplus\mathbb{C}\mathbf{v},\ \ \ \mathbf{u}\in L^{2}_{0}(\mathbb{C}/\Lambda;\mathbb{C}^{2}),\ \ \mathbf{v}\in L^{2}_{2}(\mathbb{C}/\Lambda;\mathbb{C}^{2}).$
Remark. We observe that in the case of $m(\alpha)=2$ we can write
$\ker D_{\rm S}(\alpha)=\{c\frac{\theta(z-z_{1})\theta(z+z_{1})}{\theta(z-z_{S})\theta(z+z_{S})}\mathbf{u}(\alpha),\ \ z_{1}\in\mathbb{C}/\Lambda,\ \ c\in\mathbb{C}\}.$ (3.25)
The Weierstrass identity [KhZa15, (3.3)] gives
$\theta(z-z_{S})\theta(z+z_{S})\theta(z_{1})^{2}+\theta(z_{S}-z_{1})\theta(z_{S}+z_{1})\theta(z)^{2}=\theta(z_{S})^{2}\theta(z-z_{S})\theta(z+z_{S})$
which means that, with $\mathbf{v}=F\mathbf{u}$, ($F(z)$ defined in Lemma 5)
$\frac{\theta(z-z_{1})\theta(z+z_{1})}{\theta(z-z_{S})\theta(z+z_{S})}\mathbf{u}=\frac{\theta(z_{1})^{2}}{\theta(z_{S})^{2}}\mathbf{u}+\frac{\theta(z_{S}-z_{1})\theta(z_{S}+z_{1})}{\theta(z_{S})^{2}}\mathbf{v}.$
3.5. Local structure of solutions
We start with some general comments: we consider the following non-holomorphic embedding of $\mathbb{R}^{2}\simeq\mathbb{C}$ in $\mathbb{C}^{2}$:
$F:\mathbb{C}\to\mathbb{C}^{2},\ \ \ F:z\mapsto(z,\bar{z})\in\mathbb{C}^{2}.$
The image of $F$, $X:=\{(z,\bar{z}):z\in\mathbb{C}\}$, is a totally real linear subspace (over $\mathbb{R}$) of $\mathbb{C}^{2}$ in the sense that $X\cap iX=\{0\}$. In particular, (fwe use the notation ${\rm neigh}_{\Omega}(\Omega_{0})$ for an open neighbourhood of $\Omega_{0}$ in $\Omega$, $\mathscr{O}(\Omega)$ denotes holomorphic functions in $\Omega$, $\zeta=(\zeta_{1},\zeta_{2}\in\mathbb{C}^{2}$)
$\begin{gathered}u\in\mathscr{O}({\rm neigh}_{\mathbb{C}^{2}}(X))\ \Longrightarrow\ \left\{\begin{array}[]{l}\partial_{\zeta_{1}}u(\zeta)|_{\zeta=(z,\bar{z})}=\partial_{z}[u(z,\bar{z})],\\ \partial_{\zeta_{2}}u(\zeta)|_{\zeta=(z,\bar{z})}=\partial_{\bar{z}}[u(z,\bar{z})],\end{array}\right.\\ u\in\mathscr{O}({\rm neigh}_{\mathbb{C}^{2}}(X)),\ \ u|_{X}=0\ \Longrightarrow\ u\equiv 0.\end{gathered}$ (3.26)
With this in mind, we consider the following problem (we drop $h$ or $\alpha=1/h$ which can be put into $V$):
$2D_{w}\mathbf{v}_{j}(z,w)=\begin{pmatrix}0&\alpha V(z,w)\\ \alpha&0\end{pmatrix}\mathbf{v}_{j},\ \ \ \mathbf{v}_{j}(z,\bar{z}_{0})=\mathbf{e}_{j},$ (3.27)
where $\mathbf{e}_{1}=[1,0]^{t}$, $\mathbf{e}_{2}=[0,1]^{t}$, for $z$ near $z_{0}$ and $w$ near $\bar{z}_{0}$.
Lemma 7.
The system (3.27) has unique solutions $\mathbf{v}_{j}\in\mathscr{O}({\rm{neigh}}_{\mathbb{C}^{2}}(z_{0},\bar{z}_{0});\mathbb{C}^{2})$.
Proof.
For a fixed $z$ we have the standard theory showing existence of unique solutions which are holomorphic in $w$. (In fact, we have a global solution in $w$ – see [Sj19, §7.2].) We now see that $\partial_{\bar{z}}\mathbf{u}_{j}$ solves the equation (3.27) but with the initial condition $\partial_{\bar{z}}\mathbf{u}_{j}(z,\bar{z}_{0})=0$. But that means that $w\mapsto\partial_{\bar{z}}\mathbf{u}_{j}(w,z)=0$. ∎
Lemma 8.
There exists $\mathbf{u}_{j}\in\mathscr{O}({\rm neigh}_{\mathbb{C}^{2}}(z_{0},\bar{z}_{0}))$ solving
$2D_{w}\mathbf{u}_{j}(z,w)=\begin{pmatrix}0&\alpha V(z,w)\\ \alpha&0\end{pmatrix}\mathbf{u}_{j},$
such that any solution $\mathbf{u}\in C^{\omega}({\rm neigh}_{\mathbb{C}}(z_{0}))$ solving $D_{\rm{S}}(\alpha)\mathbf{u}=0$, (see (1.2)) is given by
$\mathbf{u}(z,\bar{z})=c_{1}(z)\mathbf{u}_{1}(z,\bar{z})+c_{2}(z)\mathbf{u}_{2}(z,\bar{z}),\ \ c_{j}\in\mathscr{O}({{\rm neigh}_{\mathbb{C}}(z_{0})}).$ (3.28)
Remark. The assumption of analyticity of $V$, made throughout this paper is not necessary here. Using [KoSp57, Proposition 19.1] applied in the simplest case of $m=1$ gives the lemma for $V\in C^{\infty}(\mathbb{C})$ without the necessity of complexifying $\mathbb{C}$ to $\mathbb{C}^{2}$. Consequently, the conclusion of this section work for a more general class of potentials. We opt for a more elementary approach under the stronger assumption of real analyticity.
Proof of Lemma 8.
We have many different choices for $\mathbf{u}_{j}$ and one is given in Lemma 7. We note that $\mathbf{v}_{j}(z_{0},\bar{z}_{0})=\mathbf{e}_{j}$ and hence $\mathbf{v}_{j}(z,\bar{z})$ are linear independent for $z\in{\rm neigh}_{\mathbb{C}}(z_{0})$. If $D_{\mathrm{S}}\mathbf{u}=0$ near $z_{0}$ then $u$ is real analytic in ${\rm neigh}_{\mathbb{C}}(z_{0})\simeq{\rm neigh}_{X}(z_{0},\bar{z}_{0})$ (by the standard ellipticity result for operators with analytic coefficients). Hence it extends to $\widetilde{\mathbf{u}}\in\mathscr{{\rm neigh}}_{\mathbb{C}^{2}}(z_{0},\bar{z}_{0})$ and (as in (3.26))
$2D_{w}\widetilde{\mathbf{u}}(z,w)=\begin{pmatrix}0&\alpha V(z,w)\\ \alpha&0\end{pmatrix}\widetilde{\mathbf{u}}(z,w).$
Hence for $(z,w)\in{\rm neigh}_{\mathbb{C}^{2}}(z_{0},\bar{z}_{0})$,
$\widetilde{\mathbf{u}}(z,w)=c_{1}(z)\mathbf{u}_{1}(z,w)+c_{2}(z)\mathbf{u}_{2}(z,w),$
and, as $\widetilde{\mathbf{u}}(z,w)$ is holomorphic in both variables, the coefficients $z\mapsto c_{j}$ are also holomorphic. Restricting to $X$ gives (3.28). ∎
3.6. Rank-2 holomorphic vector bundle
The local structure of solutions shows the existence of the following rank 2 topologically trivial holomorphic vector bundle, $\mathscr{E}$, over $\mathbb{C}/\Lambda$. We have
$\mathscr{E}=\mathbb{C}/\Lambda\times\mathbb{C}^{2},\ \ \ \pi:\mathscr{E}\mapsto\mathbb{C}/\Lambda.$
To define the following covering of $\mathbb{C}/\Lambda$: for $z_{0}\in\mathbb{C}^{2}$ we choose $U={\rm neigh}_{\mathbb{C}/\Lambda}(z_{0})$ such that (3.28) holds with some linearly independent (at each $z\in U$), $\mathbf{u}_{j}$, $j=1,2$. We then define
$\begin{gathered}g_{U}=g_{U,\mathbf{u}_{1},\mathbf{u}_{2}}:U\times\mathbb{C}^{2}\to\pi^{-1}(U),\\ g_{U}(z,\zeta)\mapsto(z,\zeta_{1}\mathbf{u}_{1}(z)+\zeta_{2}\mathbf{u}_{2}(z))\subset\mathbb{C}/\Lambda\times\mathbb{C}^{2},\end{gathered}$
noting that, as $\mathbf{u}_{j}(z)$ are independent, this is a map is an isomorphism $\mathbb{C}^{2}\to\pi^{-1}(z)$. If $V$ is another neighbourhood with a local basis of the kernel of $D_{S}$ over $V$ given by $\mathbf{v}_{1},\mathbf{v}_{2}$, then Lemma 8 shows that
$\mathbf{v}_{j}(z,\bar{z})=c_{1j}(z)\mathbf{u}_{1}(z,\bar{z})+c_{2j}(z)\mathbf{u}_{2}(z,\bar{z}),\ \ \ c_{kj}\in\mathscr{O}(U\cap V).$
Hence for $z\in U\cap V$,
$g_{U}^{-1}\circ g_{V}|_{\{z\}\times\mathbb{C}^{2}}(\zeta)=C(z)\zeta,\ \ C(z):=\begin{pmatrix}c_{11}(z)&c_{12}(z)\\ c_{21}(z)&c_{22}(z)\end{pmatrix}$,
is holomorphic.
If we now include $\alpha$ we obtain a family of holomorphic vector bundles $\mathscr{E}(\alpha)$: their local sections correspond to local solutions to $D_{\rm{S}}(\alpha)\mathbf{u}=0$.
We can describe $\mathscr{E}(\alpha)$ at different values of $\alpha$ and see how it fits in the classification of holomorphic rank 2 vector bundles over tori – see [Da07, §4] for a gentle introduction to that theory.
We start with the following observation:
The determinant line bundle, $\wedge^{2}\mathscr{E}(\alpha)$, is trivial. (3.29)
Proof of (3.29).
In our choice of $\mathbf{u}_{j}(z)$, defined in $U$, we can always arrange that in $U$, $\operatorname{Wr}(\mathbf{u}_{1},\mathbf{u}_{2}):=\mathbf{u}_{1}\wedge\mathbf{u}_{2}\equiv 1$. In fact, since $D_{\bar{z}}\operatorname{Wr}(\mathbf{u}_{1},\mathbf{u}_{2})=0$, the Wronskian is holomorphic and non-vanishing, so we can replace $\mathbf{u}_{1}$ by $\mathbf{u}_{1}/\operatorname{Wr}(\mathbf{u}_{1},\mathbf{u}_{2})$. The transition matrix is given by
$C(z)=\begin{pmatrix}\ \ \mathbf{v}_{1}\wedge\mathbf{u}_{2}&\ \ \mathbf{v}_{2}\wedge\mathbf{u}_{2}\\ -\mathbf{v}_{1}\wedge\mathbf{u}_{1}&-\mathbf{v}_{2}\wedge\mathbf{u}_{1}\end{pmatrix},\ \ \ z\in U\cap V.$
But the transition factor for $\wedge^{2}\mathscr{E}(\alpha)$ is then
$\det C(z)=\det\begin{pmatrix}\mathbf{v}_{1}\wedge\mathbf{u}_{1}&\mathbf{v}_{1}\wedge\mathbf{u}_{2}\\ \mathbf{v}_{2}\wedge\mathbf{u}_{1}&\mathbf{v}_{2}\wedge\mathbf{u}_{2}\end{pmatrix}=(\mathbf{v}_{1}\wedge\mathbf{v}_{2})(\mathbf{u}_{1}\wedge\mathbf{u}_{2})\equiv 1.$
This means that constants are global sections. ∎
Case 1: $\alpha\notin{\mathcal{B}}\cup\mathcal{A}$. In this case there exists a unique global section satisfying $D_{\rm S}(\alpha)\mathbf{u}=0$, $\mathbf{u}\in C^{\omega}(\mathbb{C}/\Lambda)$. We note that $\mathbf{u}$ is a holomorphic section of $\mathscr{E}(\alpha)$ but obviously it is not a holomorphic function on $\mathbb{C}/\Lambda$. Since $\alpha\notin\mathcal{A}$, $\mathbf{u}$ does not vanish and hence it defines a trivial line subbundle, $L(\alpha)$ of $\mathscr{E}(\alpha)$. The quotient line bundle $\mathscr{E}(\alpha)/L(\alpha)$ has an explicit representation based on the proof of Proposition 1. We showed there that for $\alpha\notin{\mathcal{B}}\cup\mathcal{A}$, there exists $\mathbf{v}$ such that $D_{\rm S}(\alpha)\mathbf{v}=\mathbf{u}$. We also remark that $\mathbf{v}(z)$ is independent of $\mathbf{u}(z)$ at every point. In fact, in the notation of (3.13),
$\begin{split}2D_{\bar{z}}W(\mathbf{u},\mathbf{v})&=2D_{\bar{z}}u_{1}v_{2}+2D_{\bar{z}}v_{2}u_{1}-2D_{\bar{z}}u_{2}v_{1}-2D_{\bar{z}}v_{1}u_{2}\\ &=\alpha Vu_{2}v_{2}+(v_{1}+u_{2})u_{1}-u_{1}v_{1}-(\alpha Vv_{2}+u_{1})u_{2}=0.\end{split}$
Hence dependence at some point would imply that $W(\mathbf{u},\mathbf{v})\equiv 0$. Arguing as in (3.19), $\mathbf{u}(z,\bar{z})=f(z)\mathbf{v}(z,\bar{z})$ for some meromorphic function $f$. But then $\mathbf{u}$ would have zeros which contradicts the assumption $\alpha\notin\mathcal{A}$ and Proposition 2.
We now define the following element of $C^{\infty}(\mathbb{C};\mathbb{C}^{2})$:
$\widetilde{\mathbf{u}}(z):=\tfrac{i}{2}\bar{z}\mathbf{u}-\mathbf{v}$ (3.30)
which satisfies
$D_{\rm{S}}(\alpha)\widetilde{\mathbf{u}}=0,\ \ \ \widetilde{\mathbf{u}}(z+\gamma)=\widetilde{\mathbf{u}}(z)+\tfrac{i}{2}\bar{\gamma}\mathbf{u}(z).$ (3.31)
Since $\mathbf{u}(z)$ are $\mathbf{v}(z)$ are independent, $\mathbf{u}$ and $\widetilde{\mathbf{u}}$ can be used as a basis, near $z_{0}$ and over $\mathscr{O}({\rm neigh}_{\mathbb{C}/\Lambda}(z_{0}))$, of local solutions to $D_{\rm{S}}(\alpha)\mathbf{w}=0$.
From (3.31) we see that $\widetilde{\mathbf{u}}$ defines a non-vanishing section of $\mathscr{E}(\alpha)/L(\alpha)$. To see that $\mathscr{E}(\alpha)$ is indecomposable, suppose that $\mathscr{E}(\alpha)=L\oplus M$. Since $\mathscr{E}(\alpha)$ and $L$ are trivial so is $M$ which consequently has a global nonvanishing section, $\mathbf{w}$. Then $\mathbf{w}]=c[\widetilde{\mathbf{u}}$] in $\mathscr{E}(\alpha)/L$ and we can assume that assume $c=1$. This means that
$\mathbf{w}(z,\bar{z})=\widetilde{\mathbf{u}}(z,\bar{z})+F(z,\bar{z})\mathbf{u}(z,\bar{z}).$
Applying the equation we see that $D_{\bar{z}}F=0$, while (3.31) gives $F(z+\gamma)=F(z)-i\bar{\gamma}/2$. Consequently $f(z,\bar{z}):=F(z)+i\bar{z}/2$ is periodic and solves $\partial_{\bar{z}}f=i/2$. But that is impossible as we see by integrating both sides over $\mathbb{C}/\Lambda$.
Case 2: $\alpha\in{\mathcal{B}}$. This is the easiest case as the proof of Proposition 1 shows that $\ker D_{\rm{S}}(\alpha)$ is spanned by two independent solutions $\mathbf{u}\in L^{2}_{0}(\mathbb{C}/\Lambda;\mathbb{C})$ and $\mathbf{v}\in L^{2}_{-1}(\mathbb{C}/\Lambda;\mathbb{C})$ (see (3.12)). In particular, they never vanish and define two trivial line bundles, $L_{0}(\alpha)$ and $L_{-1}(\alpha)$ and
$E(\alpha)=L_{0}(\alpha)\oplus L_{-1}(\alpha).$
Case 3: $\alpha\in\mathcal{A}$. In view of Proposition 4, the dimension of the space of holomorphic sections of $\mathscr{E}(\alpha)$ is $m(\alpha)$, which is also the number of zeros of each section. From (3.22) we see that at each point $z$, $\{\mathbf{w}(z):\mathbf{w}\in\ker D_{\rm{S}}(\alpha)\}\simeq\mathbb{C}$, and hence $\ker D_{\rm{S}}(\alpha)$ defines a line bundle, $L(\alpha)$ over $\mathbb{C}/\Lambda$. The proof of Proposition 4 shows that this is the line bundle associated to the divisor given by the zeros of any fixed element of $\ker D_{\rm{S}}(\alpha)$.
In the notation of [Da07, Definition 2.18], $\mathscr{E}$ (we drop $\alpha$ here) is decomposable if ${\rm{Ext}}^{1}(\mathscr{E}/L,L)=\{0\}$. Since $\mathscr{E}$ is trivial, the degree of $\mathscr{E}/L$ is equal to m-m, where $m>0$ is the degree of $L$. We have, by Serre’s duality ([Da07, Proposition 2.21], with trivial $K_{X}$ for tori), ${\rm{Ext}}^{1}(\mathscr{E}/L,L)\simeq H^{1}((\mathscr{E}/L)^{*}\otimes L)\simeq H^{0}((\mathscr{E}/L)\otimes L^{*})^{*}$. But the last space is trivial as the degree of the line bundle $(\mathscr{E}/L)\otimes L^{*}$ is 2m<0-2m<0. Hence, indeed, $\mathscr{E}$ is decomposable.
As noted in (3.29), the determinant bundle of $\mathscr{E}(\alpha)$ is trivial, and if $\mathscr{E}(\alpha)=L(\alpha)\oplus M(\alpha)$ then $M(\alpha)=\det\mathscr{E}(\alpha)\otimes L(\alpha)^{*}=L(\alpha)^{*}$. Hence we can identify $\mathscr{E}(\alpha)$ with $L(\alpha)\oplus L(\alpha)^{*}$.
Figure 6. On the left: contour plot of $z\mapsto\alpha^{-1}\log|u(\alpha,z)|$, $\alpha=19.9$, where to $(2D_{\bar{z}})^{2}-\alpha^{2}U(z)U(-z))u(\alpha,z)=0$ and $U$ is the Bistritzer–MacDonald potential (2.3). We see the structure of zeros and the regions of exponential decay predicted in [HiZw25]. On the right: the plot of $z\mapsto\alpha^{-1}\log|u(\alpha,z)|$ over the the interval connected the center to the edge of the hexagon, continued through the edge and then to the translated center (top) and over the interval connecting the centers and perpendicular to the edge of the hexagon. For an animated version with varying $\alpha$ see https://math.berkeley.edu/~zworski/scalar_protected.mp4.
4. WKB structure of protected states
Numerical computations indicate that the family of protected states, $u(\alpha)$, (see Proposition 1) has a structure of a WKB state – see Figure 6 for the case of the Bistritzer–MacDonald potential (2.3) It was shown in [HiZw25] that the ($L^{2}$-normalised) protected state (in the scalar and chiral models) for that potential is exponentially small, that is of size $\exp(-c_{0}\alpha)$ in a neighbourhood of the hexagon. Numerically we see that outside of any neighbourhood of the 6 maxima of $|u(\alpha)|$ we have exponential decay of the solution. Mathematically, the key property behind the exponential decay near the hexagon is the vanishing of the Poisson bracket $\{q,\bar{q}\}|_{q^{-1}(0)}$, $q:=(2\bar{\zeta})^{2}-V(z)$ on the hexagon and its non-vanishing outside of the hexagon (except at the center). We refer to [HiZw25] for the detailed discussion of this.
The structure changes for other potentials and an interesting case is provided by
$U(z)=i\overline{U_{BM}(z)}^{2},$ (4.1)
which satisfies the conditions (2.2) but for which $\{q,\bar{q}\}_{q^{-1}(0)}\equiv 0$. That implies an existence of a global solution to the eikonal equation and suggests different WKB structure.
In this section we will first discuss the zero creation at the corners of the hexagon for the Bistritzer–MacDonald potential. We will then present a heuristic derivation of a WKB state for the model based on (4.1) and an argument towards the proof of Theorem 2.
4.1. Mechanism for the creation of zeros for the Bistritzer–MacDonald potential
For the scalar model numerical calculation show that the real elements of $\mathcal{A}$ have multiplicity 2. In view of the arguments presented in §3.4 this means that the protected state $v(\alpha)$, $D_{\rm{S}}(\alpha)v(\alpha)=0$ has simple zeros (in the sense of (3.20)) at the corners of the hexagon. Using (1.2) we see that
$v(\alpha,z)=\begin{pmatrix}\alpha^{-1}2D_{\bar{z}}u(\alpha,z)\\ u(\alpha,z)\end{pmatrix},\ \ \ ((2D_{\bar{z}})^{2}-\alpha^{2}V(z))u(\alpha,z)=0.$
The structure of $u(\alpha,z)$ was illustrated in Figure 6.
In this section we will describe an approximate local mechanism for the creation of zeros of $u$ and $\partial_{\bar{z}}u$ at the corner. The symmetries of $u$ will play a crucial role here.
We start by describing $V(z)=U_{\rm{BM}}(z)U_{\rm{BM}}(-z)$, where $U_{\rm{BM}}$ is given in (2.3) near the corner of the hexagon $z_{S}$:
$V(z_{S}+w)=-ia\bar{w}+bw^{2}+\mathcal{O}(|w|^{3}),\ \ \ a=\tfrac{32}{3}\pi^{3},\ \ b=\tfrac{32}{9}\pi^{4}+\mathcal{O}(|w|^{3}).$ (4.2)
The model operator near the corner (writing $z=w$) is then given by
$P=(2D_{\bar{z}})^{2}+\alpha^{2}(ia\bar{z}-bz^{2}).$ (4.3)
To represent the solutions using Airy functions we write
$z=-i\beta w,\ \ \ \bar{z}=i\beta\bar{w},\ \ D_{\bar{z}}=-\beta^{-1}\partial_{\bar{w}},$
so that if $\beta=(2/\alpha)^{\frac{2}{3}}a^{-\frac{1}{3}}$ then
$\begin{split}P&=(2h/\alpha)^{2}\partial_{\bar{w}}^{2}-a\alpha\bar{w}+b\alpha^{2}w^{2}\\ &=4(\alpha/2)^{4/3}a^{\frac{1}{3}}(\partial_{\bar{w}}^{2}-\bar{w}+c\alpha^{-\frac{2}{3}}w^{2}),\ \ \ c:=2^{\frac{2}{3}}b/a^{\frac{4}{3}}=\tfrac{2}{3}(\tfrac{16}{3})^{\frac{1}{3}}>0.\end{split}$
We keep in mind that
$w=i\beta(\alpha)z,\ \ \ \beta(\alpha):=(2/a)^{\frac{2}{3}}a^{-\frac{1}{3}}.$ (4.4)
Let $u$ denote the solution in the original variables of (1.1). We want to deduce the symmetries of (local) solutions to $Pv=0$ pretending that the operators agree. In view of (3.5) and the fact that $\omega z_{S}\equiv z_{S}\!\!\mod\!\Lambda$, $\bar{z}_{S}=-z_{S}\!\!\mod\!\Lambda$, we have
$\begin{gathered}v(w)=u(z_{S}+\omega i\beta w)=u(\omega(z_{S}+i\beta w))=u(z_{S}+\omega i\beta w)=v(\omega w),\\ v(\bar{w})=u(z_{S}+i\beta\bar{w})=u(\overline{-z_{S}-i\beta w})=\overline{u(-z_{S}-i\beta w)}=\overline{u(z_{S}+i\beta w)}=\overline{v(w)}.\end{gathered}$ (4.5)
(Note that this last consequence is global in the sense that we use global properties of our solution.)
Figure 7. Illustration of $g_{0}(0,\alpha)$ in (4.8) and (4.9) – it is calculated by numerically evaluating the Wronskian of the protected state with $Ai_{2}(\bar{w}-\alpha^{-\frac{2}{3}}cw^{2})$. We see exponential decay explained in [HiZw25, Theorem 1] and the zeros at the magic angles of the scalar model for the potential (2.3).
We recall some basic facts about Airy functions [^2] See https://mtaylor.web.unc.edu/wp-content/uploads/sites/16915/2018/04/airyf.pdf for a quick but thorough introduction.. The first Airy function is given by
$Ai(s)=\frac{1}{2\pi}\int_{\mathbb{R}}e^{ist+it^{3}/3}dt,\ \ \ \partial_{s}^{2}Ai(s)-sAi(s)=0.$
By either considering the equation satisfied by $Ai$ or by deforming the contour in the integral we see that $\mathbb{C}\ni\xi\mapsto Ai(\xi)$ is holomorphic and $Ai(\overline{\xi})=\overline{Ai(\xi)}$. That allows to define
$Ai_{j}(\xi):=\sum_{k=0}^{j}\omega^{kj}Ai(\omega^{k}\xi).$ (4.6)
It follows that $Ai_{1}\equiv 0$ and that $Ai_{0}$ and $Ai_{2}$ give a basis of solutions to the Airy equations ($Ai_{0}(0)=3Ai(0)\neq 0$, $Ai_{0}^{\prime}(0)=0$, $Ai_{2}(0)=0$, $Ai_{2}^{\prime}(0)=3Ai^{\prime}(0)\neq 0$). We also have the following properties
$\overline{Ai_{j}(\xi)}=Ai_{j}(\overline{\xi}),\ \ \ Ai_{j}(\omega\xi)=\omega^{-j}Ai_{j}(\xi).$ (4.7)
Figure 8. On the left a snapshot from the movie referenced in Figure 6 showing the evolutions of zeros at the corner; on the right a snapshot from https://math.berkeley.edu/~zworski/airy_movie.mp4 showing the same structure for (4.8) with $g_{0}$ given by (4.10) with $c_{j}=1$, $j<4$, $c_{4}=0$.
We now consider solutions to $Pv=0$ where $P$ is (4.3) but in coordinates (4.4). In view of (4.2) these solutions approximate solutions to $P(\alpha,0)u=0$ near the corners of the hexagon. They are given by
$v(w,\bar{w})=f_{0}(w,\alpha)A_{0}(\bar{w}-\alpha^{-\frac{2}{3}}cw^{2})+f_{2}(w,\alpha)A_{2}(\bar{w}-\alpha^{-\frac{2}{3}}cw^{2}),$ (4.8)
where $f_{j}$, $j=0,2$ are holomorphic in $w$. Combining (4.5) with (4.7) we obtain, with $g_{j}$, $j=1,2$, holomorphic,
$f_{0}(w,\alpha)=g_{0}(w^{3},\alpha),\ \ \ f_{2}(w,\alpha)=wg_{1}(w^{3},\alpha),\ \ \ g_{j}(\bar{\zeta},\alpha)=\overline{g_{j}(\zeta,\alpha)}.$ (4.9)
Numerically we can compute $g_{j}$’s by taking Wronskians of the actual solution to $P(\alpha,0)u=0$ with $Ai_{j}(\bar{w}-\alpha^{-\frac{2}{3}}cw^{2})$. The resulting functions $g_{j}$ are close to being holomorphic (the Airy functions provide approximate solutions for $P(\alpha,0)$ – see (4.2); they are holomorphic when solutions to $Pv=0$ are decomposed using (4.8)) and satisfy
$\begin{gathered}g_{0}^{\prime}(0,\alpha)\neq 0,\ \ \ g_{2}(0,\alpha)\neq 0.\end{gathered}$
and the zeros of $g_{0}(0,\alpha)$ determine the distribution of $\alpha$’s. Figure 7 shows the plot of the numerically computed $\log|g_{0}(0,\alpha)|$ as function of $\alpha$ (using (4.8) and (4.9)). Exponential decay and the way zeros are created suggests that
$g_{0}(\zeta,\alpha)\simeq e^{-c_{0}\alpha}(c_{1}\zeta+c_{2}\sin(c_{3}\alpha+c_{4})).$ (4.10)
That choice of $g_{0}$ in (4.8) and (4.9) matches the the observed mechanism of zero creation – see Figure 8.
Figure 9. An attempt to illustrate the characteristic variety $q^{-1}(0)$ for the classical symbol of (1.1), $q(x,\xi)=(2\bar{\zeta})^{2}-V(z)$. It is a genus two surface with singular projection to the base torus above the vertices of the hexagon (the stacking points $\pm z_{S}$) and a single singularity above the center of the hexagon. That singularity is almost a normal crossing $\bar{\zeta}^{2}-z^{2}=0$ but the two surfaces, $\bar{\zeta}\simeq\pm z$ crossing transversally at the center (and momentum zero) are not smooth but $C^{2,\gamma}$, $\gamma<1$. Elsewhere, the surfaces is real analytic.
Ideally one would like to connect solutions approximated by (4.8) with other local WKB solutions to obtain an approximation for $g_{0}(0,\alpha)$ coming from a global solution: knowing $c_{3}$ in (4.10) would then given a spacing in the values of real $\alpha$’s as in (1.7). An attempt in that direction was made in [Re*21] in a somewhat different spirit than proposed in this section. The difficulty lies in a complicated structure of the characteristic variety $q^{-1}(0)$, $q(x,\xi)=(2\bar{\zeta})^{2}-V(z)$, $\zeta=\frac{1}{2}(\xi_{1}-i\xi_{2})$, $z=x_{1}+ix_{2}$ (this is the classical observable which is semiclassical quantised, with $h=1/\alpha$ to (1.1) – see [HiZw25]). An attempt to illustrate that structure is made in Figure 9. The exponential decay of the solution near the hexagon adds to the many complications.
Figure 10. The characteristic variety $q^{-1}(0)$ for the classical symbol of (1.1), $q(x,\xi)=(2\bar{\zeta})^{2}-V(z)$ with V(z)=U(z)U(z)V(z)=U(z)U(-z) and $U(z)=i\overline{U_{\rm{BM}}(z)}^{2}$. It is a union of graphs of $\pm d\varphi(z)$, where $\varphi$ is real and is given in (4.13). These two real Lagrangian tori touch at the edges of the hexagon and at its center. That complicates the WKB structure of solutions which nevertheless is much simpler than that for the characteristic variety shown in Figure 9.
4.2. An example with a quantisation condition for magic alphas
We now considerUU given by (4.1). The properties (2.2) are inherited from the same properties being satisfied by $U_{\rm{BM}}$. Hence, the set of magic alphas, $\mathcal{A}$ is well defined both for the corresponding chiral and scalar models. Here we will consider the latter. What changes is the structure of the characteristic variety $q^{-1}(0)$, $q:=(2\bar{\zeta})^{2}-V(z)$. It consists of two Lagrangian tori with three high tangency common points – it is shown in Figure 10.
The simpler structure of the characteristic variety is due the existence of a global phase function. To see it we write the potential as
$\begin{gathered}V(z)=W(z)^{2}=U(z)U(-z),\ \ \ U(z):=i\overline{U_{\rm{BM}}(z)}^{2},\\ W(z):=i\overline{U(z)U(-z)}=-2i\left(\frac{4\pi}{3}\right)^{2}\sum_{l=0}^{2}\omega^{l}\cos\left(\left\langle z,\omega^{l}{4\pi i}/{\sqrt{3}}\right\rangle\right).\end{gathered}$ (4.11)
The eikonal equation for $(2D_{\bar{z}})-\alpha^{2}W(z)^{2}$ (where we think of $\alpha=1/h$ as a semiclassical parameter) reads
$2\partial_{\bar{z}}\varphi(z)=W(z),$ (4.12)
and it is solved by a real valued function:
$\varphi(z):=\frac{8\pi}{3\sqrt{3}}\sum_{l=0}^{2}\sin\left(\left\langle z,{4\pi i}/{\sqrt{3}}\omega^{l}\right\rangle\right).$ (4.13)
Hence
$\begin{gathered}q^{-1}(0)=\{(x,d_{x}\varphi(x)):x\in\mathbb{C}/\Lambda\}\cup\{(x,-d_{x}\varphi(x)):x\in\mathbb{C}/\Lambda\},\\ q(x,\xi)=(2\bar{\zeta})^{2}-W(z)^{2},\ \ \ \zeta=\tfrac{1}{2}(\xi_{1}-i\xi_{2}),\ \ z=x_{1}+ix_{2}.\end{gathered}$ (4.14)
In particular, we have
$\{q,\bar{q}\}|_{q^{-1}(0)}=0,$ (4.15)
and consequently there is no mechanism for exponential decay of solutions to $P(\alpha,0)u=0$ as in the case of $U=U_{\rm{BM}}$ – see [HiZw25].
The phase function (4.13) enjoys the following natural symmetries:
$\begin{gathered}\varphi(-z)=\varphi(\bar{z})=-\varphi(z),\ \ \ \varphi(\tau z)=-\varphi(z),\ \ \ \tau=e^{-i\pi/3}.\end{gathered}$ (4.16)
For future reference we also consider the behaviour of $\varphi$ at the vertex, $z=z_{S}$ and at the center $z=0$, of the hexagon. For that we first note that (4.2) gives
$W(z_{S}+w)=i\overline{V_{\rm{BM}}(z)}=i(-iaw+b\bar{w}^{2}+\mathcal{O}(|w|^{3})))=aw+\mathcal{O}(|w|^{2}),$
so that
$\varphi(z_{S}+w)=\tfrac{1}{2}a|w|^{2}+\mathcal{O}(|w|^{3}),\ \ \ a=\tfrac{32}{3}\pi^{3}.$ (4.17)
At the center we have
$\begin{split}\varphi(z)&=\frac{8\pi}{3\sqrt{3}}\sum_{\ell=0}^{2}\sin\left(\frac{2\pi i}{\sqrt{3}}(\bar{z}\omega^{\ell}-z\bar{\omega}^{\ell})\right)=-i\frac{32\pi^{4}}{81}\sum_{\ell=0}^{2}\left(\bar{z}\omega^{\ell}-z\bar{\omega}^{\ell}\right)^{3}+\mathcal{O}(|z|^{5})\\ &=-i\frac{32\pi^{4}}{27}(\bar{z}^{3}-z^{3})+\mathcal{O}(|z|^{5})=-\frac{64\pi^{4}}{27}\operatorname{Im}z^{3}+\mathcal{O}(|z|^{5}).\end{split}$ (4.18)
Figure 11. The protected state for the potential (4.1) over a segment on the imaginary axis (it is real valued there) compared to WKB approximations near the center of the hexagon (based on Hankel functions) and away from the center (based on a standard WKB approximation multiplied by a suitable holomorphic function). Although we do not present a rigorous proof of this approximation the agreement with the numerics is quite striking. For an animated version with varying $\alpha$, see https://math.berkeley.edu/~zworski/on_the_edge.mp4.
We will now propose an approximate solution to $P(\alpha,0)u=0$ satisfying the symmetries of the protected state (3.5). We will construct $u$ over the closed parallelogram, $\Sigma$, spanned by $\omega$ and $\omega^{2}$ (this points are $\Lambda$-congruent to 0, the centre of the hexagon) and $(\omega-\omega^{2})/3$, $2(\omega-\omega^{2})/3$ (vertices of the hexagon). We remark that
$\Sigma_{\pm}:=\Sigma\cap\{\pm\operatorname{Re}z\geq 0\},\ \ \ \Sigma_{0}:=\Sigma\cap\{\operatorname{Re}z=0\},$ (4.19)
where the $\pm$ sets are fundamental domains for the actions of the lattice $\Lambda$ and $z\mapsto e^{i\pi/3}z$ while $\Sigma_{0}$ is the edge of the hexagon. The approximate solution is pieced from an approximate solution valid away from $\pm\frac{1}{2}$ (”near the edge”) and an approximate solution valid near $\pm 12$ (”near the centre”).
Figure 12. Bottom: the (dominant) real part of the protected state for $\alpha=1$ over the fundamental domain given by the hexagon. Top: the matched ansätze (4.20) and (4.25). The strong agreement in the “eyeball norm” suggests a rigorous agreement of finer approximations – see Figure 13 for some numerical comparisons. A rotating version of this figure can be found at
https://math.berkeley.edu/~zworski/rotate_compare.mp4.
The ultimate goal is to count zeros on the edge as a function of $\alpha$: a creation of a zero at the vertex at a magic $\alpha$ produces two zeros on the edge (see Figure 6). Hence counting of those zeros will determine the distribution of real magic $\alpha$’s.
Figure 13. Numerical comparison of the rough WKB approximation and the exact solution. The global relative error is poor due to the quality of the matching; the error for the Hankel function approximation ($0<\operatorname{Im}z<0.15$) is also not that good as it is based on the Taylor expansion near 0; on the other hand the approximation on the edge of the hexagon ($\operatorname{Im}z>\sqrt{3}/3$) is excellent. Numerically we also see that the imaginary part is small.
4.2.1. Approximate solution near the edge
From (3.5) we see that we seek approximate solution which is real on the boundary of $\Sigma_{\pm}$. The protected state is not real everywhere but numerically we see that the imaginary part decays with $\alpha$. We expect that this is related to the structure of the characteristic variety (4.14), (4.15) but at the moment we do not have a rigorous argument for that.
Leading terms in a WKB solutions to $P(\alpha,0)u=0$ are given by $W(z)^{-\frac{1}{2}}\exp(\pm i\varphi(z))$. The approximation of the protected state should be real and bounded on $\Sigma_{0}$ (the edge of the hexagon). Since we can multiply the approximate solutions by (local) holomorphic functions this suggests the following ansatz in $\Sigma$, near $\Sigma_{0}\cap\{\operatorname{Re}z=0\}$:
$u_{\rm E}(z)=(W_{\rm{hol}}(z)/W(z))^{\frac{1}{2}}\cos(\alpha\varphi(z)),\ \ \ W_{\rm{hol}}\in\mathscr{O}(\Sigma),\ \ \ W_{\rm{hol}}|_{\Sigma_{0}}=W|_{\Sigma_{0}}.$ (4.20)
In words, $W_{\rm{hol}}$ is the holomorphic extension of the real analytic function $W(z)$ from $\Sigma_{0}$ to $\Sigma$:
$W_{\rm{hol}}(z)=-2i\left(\frac{4\pi}{3}\right)^{2}\left(\cos\left(4\pi iz/{\sqrt{3}}\right)-\cos\left(2\pi iz/{\sqrt{3}}\right)\right),$
The ansatz (4.20) is in remarkable with the protected state away from the $\pm\frac{1}{2}$ that is away from the center of the hexagon – see Figure 12. That is particularly striking on the edge continued (in a different fundamental domain) to the center – see Figure 11. We stress that a much finer analysis would be needed to justify this agreement. That has to involve matching the ansatz (4.20) with the rougher ansatz near the center (4.25).
4.2.2. Approximate solution near the centre
Near the center we will use a simplified version of $\varphi(z)$ obtained by using the leading term in (4.18):
$\varphi_{0}(z):=ai(z^{3}-\bar{z}^{3}),\ \ \ W_{0}(z):=2\partial_{\bar{z}}\varphi_{0}(z)=-6ia\bar{z}^{2},\ \ \ a:=\frac{32\pi^{4}}{27}.$ (4.21)
(Eventually, when matching, we will use the lattice to move 0 to the $\omega,-\omega^{2}$ vertices of $\Sigma$ and vice versa.) To understand the structure of solutions to $((2D_{\bar{z}})^{2}-\alpha^{2}W_{0}(z)^{2})u(z)=0$, we note that the leading term of a WKB solution is given by
$u_{\pm}(\bar{z}):=W_{0}(z)^{-\frac{1}{2}}e^{\pm i\varphi_{0}(z)}=c_{0}\bar{z}^{-1}e^{\pm\alpha a\bar{z}^{3}}e^{\mp\alpha az^{3}}.$ (4.22)
Since we are free to multiply the approximate local solution by holomorphic functions, this suggests a substitution
$u(\bar{z})=\bar{z}^{-1}f(i\alpha a\bar{z}^{3})\ \Longrightarrow\ w^{2}f^{\prime\prime}(w)+(w^{2}+\tfrac{1}{9})f(w)=0,\ \ \ w:=i\alpha a\bar{z}^{3}.$ (4.23)
The equation for $f$ is essentially the Bessel equation: if $f(w)=w^{1/2}g(w)$, then then
$w^{2}g^{\prime\prime}(w)+wg^{\prime}(w)+(w^{2}-\tfrac{1}{36})g(w)=0.$
This is the Bessel equation and in our case (in view of (4.22)) is it natural to consider solutions as linear combinations of Hankel functions (see [AbSt83, §9.1] for definitions and properties of these functions): $g(w)=H^{(j)}_{1/6}(w)$, $j=1,2$,
$\begin{gathered}H^{(j)}_{1/6}(w)=w^{1/6}F_{1}(w^{2})-(-1)^{j}iw^{-1/6}F_{2}(w^{2}),\ \ F_{j}(0)\neq 0,\ \ F_{j}(w)=\overline{F_{j}(\bar{w})},\\ H^{(j)}_{1/6}(w)\simeq(2/\pi)^{\frac{1}{2}}w^{-\frac{1}{2}}e^{-(-1)^{j}i(w-\frac{\pi}{3})},\ \ |w|\to\infty,\ \ -\pi<\arg w<2\pi.\end{gathered}$ (4.24)
(We define the branch of $w^{\frac{1}{6}}$ for $\pi<\arg w<2\pi$ and the values on the negative axis correspond to $\arg w=\pi$.) Returning to (4.23), we see that the two standard solutions to $(2D_{\bar{z}})^{2}-\alpha^{2}W_{0}(z))u=0$ are
$u_{j}(\bar{z})=\bar{z}^{1/2}H_{1/6}^{(j)}\left(ia\alpha\bar{z}^{3}\right),\quad j=1,2,\ \ \ a=32\pi^{4}/{27}.$
We note that in view of (4.24), $\bar{z}\mapsto u_{j}(\bar{z})$ are entire (once we choose the correct branch of $\bar{z}\mapsto\bar{z}^{1/2}$. We also note that in view of (4.24) the complex conjugates, $\overline{u_{j}(z)}$ also solve the equation.
We use $u_{j}(\bar{z})$ as approximate solutions in
$\Omega_{\pm}^{\delta}:=\pm\{|z|<\delta,|\arg z|\leq\pi/6\},\ \ \ \Omega_{+}^{\delta}+\omega={\rm neigh}_{\Sigma}(\omega),\ \ \Omega_{-}^{\delta}-\omega^{2}={\rm neigh}_{\Sigma}(-\omega^{2}),$
that is near the left and right corners of $\Sigma$. That is the region where (4.20) fails to provide an approximation.
We need to take a combination of $u_{j}(\bar{z})$, $\overline{u_{j}(z)}$ with holomorphic coefficients which is smooth at 0 when extended from $\Omega_{+}^{\delta}$ to a neighbourhood of 0 using the symmetry $z\mapsto e^{i\pi/3}$ and matches the ansatz (4.20). To satisfy these conditions we arrived at the following solution to $((2D_{\bar{z}})^{2}-\alpha^{2}W_{0}(z))u=0$:
$u_{\rm{C}}(z)=i|z|\left(H^{(1)}_{1/6}(ai\alpha\bar{z}^{3})H^{(2)}_{1/6}(i\alpha az^{3})-\overline{H^{(2)}_{1/6}(ai\alpha\bar{z}^{3})}\,\overline{H^{(2)}_{1/6}(i\alpha az^{3})}\right).$ (4.25)
We consider $u_{\rm C}$ as an approximate solution to $((2D_{\bar{z}})^{2}-\alpha^{2}W(z))u=0$ near the centre. A finer ansatz involving a more complicated equation than (4.23), depending on $z$ as a parameter, would be needed to produce a rigorous agreement. Nevertheless, the agreement shown in Figures 11 and 12 is remarkable.
4.2.3. Quantisation rule for magic alphas
We expect real $\alpha\in\mathcal{A}$ to have multiplicity two and the discussion in §3.4 then shows that at such $\alpha$’s, $u(\alpha)$ and $D_{\bar{z}}u(\alpha)$ of Proposition 1 has simple zeros at the corners of the hexagon (in the sense of (3.20) and the discussion following it). We could use the ansatz (4.20) to find $\alpha$’s for which $u_{\rm C}(\alpha,i/\sqrt{3})=0$ and that provides a good agreement (in view of the very good agreement with $u$ on the edge of the hexagon – see Figure 11.
A more robust way, potentially applicable in other cases such as the scalar model based on the Bistritzer–MacDonald potential, is given by counting zeros of $u(\alpha)$ (the family of protected states in Theorem 1) on the edge of the hexagon.
“Proof” of Theorem 2.
Every time a zero occurs at the corner two zeros are added to the zeros on the edge – see the movies references in Figures 6 (for the Bistritzer–MacDonald potential) and 11 (for the potential (4.1)). Hence of that number is given by $N(\alpha)$ then the spacing between magic alpha’s, $\Delta\alpha$ is given by solving $N(\alpha+\Delta\alpha)-N(\alpha)=2$. The approximation of $u(\alpha)$ given by (4.20) is given by $\cos(\alpha\varphi(z))$. On the edge we have
$\varphi(z(t))=\frac{8\pi}{3\sqrt{3}}\left(\sin\left(\frac{4\pi}{3}t\right)-2\sin\left(\frac{2\pi}{3}t\right)\right),\ \ z(t)=it/\sqrt{3},\ \ \ 1\leq t\leq 2,$
from which we see that $t\mapsto\varphi(z(t))$ is strictly increasing. Hence $N(\alpha)\simeq\alpha(\varphi(2)-\varphi(1))$ and
$\Delta\alpha\simeq\frac{2}{\varphi(2)-\varphi(1)}=\frac{1}{4}.$
This provides a heuristic argument for (1.9). ∎
5. A toy model
A much simpler version of (1.1) is given when $V(z)=W(z)^{2}$ and $W\neq 0$. In that case the characteristic variety $q^{-1}(0)$, $q=(2\bar{\zeta})^{2}-W(z)^{2}$ is given by a union of two disjoint tori $\{(x,\xi):2\bar{\zeta}=\pm W(z)\}$, $\zeta=\frac{1}{2}(\xi_{1}-i\xi_{2})$. These tori are Lagrangian if and only of $\operatorname{Im}\partial_{z}W(z)\equiv 0$ but we are interested in the general case of complex trigonometric polynomials $W(z)$ satisfying the condition
$W_{0}:=|\mathbb{C}/\Lambda|^{-1}\int_{\mathbb{C}/\Lambda}W(z)dm(z)\neq 0.$ (5.1)
Here $\Lambda$ is a lattice of periodicity of $W$ (not necessarily the lattice we considered so far).
Figure 14. The characteristic variety $q^{-1}(0)$ for the classical symbol of (1.1), $q(x,\xi)=(2\bar{\zeta})^{2}-V(z)$ with $V(z)=W(z)^{2}$ and $W(z)\neq 0$. It is a union of two disjoint tori projecting diffeormorphically to the base torus. The tori are Lagrangian only if $\operatorname{Im}\partial_{z}W(z)\equiv 0$. Flat bands do not occur for this model but we can ask for the values of $\alpha$ for which $\ker_{H^{2}(\mathbb{C}/\Lambda)}P(\alpha,k)\neq\{0\}$. An answer in a special case is given in Theorem 3.
We recall that for any $V\in L^{\infty}(\mathbb{C}/\Lambda;\mathbb{C})$, $k\in\mathbb{C}\setminus\Lambda^{*}$, $\alpha\mapsto P(\alpha,k)^{-1}$ is a meromorphic family of operators. In fact, For $k\notin\Lambda^{*}$, $(2D_{\bar{z}}+k)^{-2}:L^{2}(\mathbb{C}/\Lambda)\to H^{2}(\mathbb{C}/\Lambda)$ exists and hence
$P(\alpha,k)=(2D_{\bar{z}}+k)(I-\alpha^{2}(2D_{\bar{z}}+k)^{-2}V(z).$
Since $\alpha\mapsto\alpha^{2}(2D_{\bar{z}}+k)^{-2}V(z)$ is a holomorphic family of compact operators, meromorphic Fredholm theory (see [DyZw19, Theorem C.8]) gives the meromorphy of $\alpha\mapsto(I+\alpha^{2}(2D_{\bar{z}}+k)^{-2}V)^{-1}$ and hence of $P(\alpha,k)^{-1}$ (we have the required invertibility at one point by putting $\alpha=0$).
Under an additional assumption it is easy to that there are no protected states at $k\in\Lambda^{*}$:
Lemma 9.
Suppose that $V$ satisfies
$\int_{\mathbb{C}/\Lambda}V(z)dm(z)\neq 0.$ (5.2)
Then, in the notation of (1.1), and for all $k\in\mathbb{C}$, $\alpha\mapsto P(\alpha,k)^{-1}$ is a meromorphic family of operators. In particular $\ker P(\alpha,0)=\{0\}$ except for a discrete set of $\alpha$’s.
Proof.
For $k\in\Lambda^{*}$ we can restrict ourselves to the case of $k=0$ (conjugation by $e^{i\langle k,z\rangle}$ reduces to this case). Let us assume for simplicity that $|\mathbb{C}/\Lambda|=1$. We have the following Grushin problem (see [DyZw19, §C.1]) for $P(0,0)$:
$\begin{gathered}\begin{pmatrix}(2D_{\bar{z}})^{2}&R_{-}\\ R_{+}&0\end{pmatrix}^{-1}:H^{1}(\mathbb{C}/\Lambda;\mathbb{C})\oplus\mathbb{C}\to L^{2}(\mathbb{C}/\Lambda;\mathbb{C})\oplus\mathbb{C},\\ R_{-}u_{-}=u_{-}\ \ \ R_{+}u=\langle u,1\rangle.\end{gathered}$
The inverse is given by
$\begin{pmatrix}E&E_{+}\\ E_{-}&0\end{pmatrix},\ \ \ E_{+}=R_{-},\ \ E_{-}=R_{+},\ \ Eu=(2D_{\bar{z}})^{-2}(v-\langle v,1\rangle).$
We now consider $P(\alpha,0)$ as a perturbation of $P(0,0)$: for small values of $\alpha$ the Grushin problem with the same $R_{\pm}$ remains invertible with the inverse (see [DyZw19, C.1.7]):
$\begin{pmatrix}E(\alpha)&E_{+}(\alpha)\\ E_{-}(\alpha)&E_{-+}(\alpha)\end{pmatrix},\ \ E_{-+}(\alpha)=\sum_{k=1}^{\infty}(-1)^{k}\alpha^{2k}E_{-}V(EV)^{k-1}E_{+}.$ (5.3)
We recall from [DyZw19, (C.1.1)] that $P(\alpha,0)$ is invertible if and only if $E_{-+}(\alpha)\neq 0$. The leading term in the expansion of $E_{-+}(\alpha)$ is given by $\alpha^{2}\int_{\mathbb{C}/\Lambda}V(z)dm(z)$. Hence (5.2) implies that $P(\alpha,0)^{-1}$ exists for $\alpha\in D(0,\varepsilon)\setminus\{0\}$. Since $\alpha\mapsto P(\alpha,0)$ is a holomorphic family of Fredholm operators we can again apply analytic Fredholm theory to obtain global meromorphy of $P(\alpha,0)$. ∎
Remark. Suppose that $\Gamma\subset\Lambda^{*}$ such that $0\notin\Gamma$ and for $\gamma,\gamma^{\prime}\in\Gamma$, $\gamma+\gamma^{\prime}\in\Gamma$. Suppose that
$V(z)=\sum_{\gamma\in\Gamma}a_{\gamma}e^{i\langle z,\gamma\rangle}\in C^{\infty}(\Lambda/\Gamma),\ \ \ a_{\gamma}=\mathcal{O}(|\gamma|^{-\infty}).$ (5.4)
Then $\ker P(\alpha,0)\neq\{0\}$ for all $\alpha$’s. (Of course the condition (5.2) is obviously violated). This follows from (5.3). In fact, if $C^{\infty}_{\Gamma}$ denotes the set of $V^{\prime}s$ of the form (5.4) then
$E:C^{\infty}_{\Gamma}\to C^{\infty}_{\Gamma},\ \ \ V:C^{\infty}_{\Gamma}\to C^{\infty}_{\Gamma},\ \ \ E_{-}:C^{\infty}_{\Gamma}\to 0.$
(Here as above $V$ as an operator denotes multiplication by $V$.)
In view of the lemma it natural to ask about the set
$\mathcal{A}(k):=\{\alpha:\ker P(\alpha,k)\neq\{0\}\}.$ (5.5)
with the multiplicity of element of $\mathcal{A}(k)$, $m(\alpha,k)$, defined by
$m(\alpha,k):=\frac{1}{2\pi i}\operatorname{tr}_{L^{2}(\mathbb{C}/\Lambda)}\oint_{\alpha}P(\zeta,k)^{-1}\partial_{\zeta}P(\zeta,k)d\zeta,$ (5.6)
where the integral is over a positively oriented circle including $\alpha$ as the only possible pole of $\zeta\mapsto P(\zeta,k)^{-1}$. If $k\in\Lambda^{*}$ this assumes meromorphy of $\zeta\mapsto P(\zeta,k)^{-1}$.
In this section we will consider this question for the simplest case $\Lambda=\mathbb{Z}^{2}$ and $W(x,y)=W(x)$, $(x,y)\in\mathbb{R}^{2}$. That amounts to the study of the operator (1.13). We present a self-contained argument which indicates that even in the simplified setting (compared to that in §4 a rigorous derivation becomes quite involved.
5.1. Semiclassical formulation
We consider the differential operator
$P(\alpha,h)=(hD_{z})^{2}-\alpha^{2}W(z)^{2},\quad z=x+iy\in\mathbb{C},\quad D_{z}=\frac{1}{2i}(\partial_{x}-i\partial_{y})$
where $\alpha\in\mathbb{C}$, $h>0$, and $W$ is an entire function. We will study the limit $h\to 0$ and assume that
$C_{0}^{-1}\leq|\alpha|\leq C_{0}$ (5.7)
for some constant $C_{0}$. We should the problem scales and we could consider $|\alpha|\gg 1$ instead of (5.7) and $h\to 0$. It is however natural in the context of WKB to use real $h\to 0$ as the asymptotic parameter.
We study holomorphic solutions $u(z)$ to the equation
$P(\alpha,h)u=0.$ (5.8)
We first review some basic properties of second order holomorphic ODEs. For a holomorphic function $u$, define the vector-valued holomorphic function
$\mathbf{u}(z)=(u(z),hD_{z}u(z))\ \in\ \mathbb{C}^{2}.$ (5.9)
The Cauchy problem for the equation (5.8) is well-posed: for any $z_{0}\in\mathbb{C}$, the map $u\mapsto\mathbf{u}(z_{0})$ is an isomorphism from the space of solutions to (5.8) onto $\mathbb{C}^{2}$ – see for instance [Sj19, §7.2].
For $z_{0},z_{1}\in\mathbb{C}$, define the transition matrix $\mathcal{M}_{z_{0}\to z_{1}}:\mathbb{C}^{2}\to\mathbb{C}^{2}$ as follows: for each solution $u$ to (5.8) we have
$\mathbf{u}(z_{1})=\mathcal{M}_{z_{1}\leftarrow z_{0}}\mathbf{u}(z_{0}),\ \ \ \mathcal{M}_{z_{1}\leftarrow z_{0}}=\mathcal{M}_{z_{1}\leftarrow z_{0}}(\alpha,h).$ (5.10)
For two holomorphic functions $u,v$, define their Wronskian (it differs from the Wronskian used in §3 by a factor of $h$):
$\operatorname{Wr}(u,v):=u(hD_{z}v)-(hD_{z}u)v.$ (5.11)
We will use the identity
$hD_{z}\operatorname{Wr}(u,v)=uP(\alpha,h)v-vP(\alpha,h)u.$ (5.12)
If $u,v$ are both solutions to (5.8), then (5.12) shows that $\operatorname{Wr}(u,v)$ is constant. This implies that
$\det\mathcal{M}_{z_{1}\leftarrow z_{0}}=1.$ (5.13)
Finally, let $\Phi(z)$ be an antiderivative of $W$:
$\partial_{z}\Phi=W.$ (5.14)
5.2. WKB approximation
We will now study the solutions to (5.8) using the global phase function
5.2.1. A growth bound
From now on in this section, we fix a smooth non-self-intersecting curve:
$\gamma:[0,1]\to\mathbb{C}\setminus W^{-1}(0),\quad\gamma^{\prime}\neq 0,$ (5.15)
We henceforth denote by $C_{\gamma}$ a constant depending only on $\gamma$, $W$, and the constant $C_{0}$ in (5.7), whose precise meaning might change from place to place.
Define the real-valued function
$\psi_{\gamma,\alpha}(t):=\operatorname{Im}\big{(}\alpha\Phi(\gamma(t))\big{)},\quad t\in[0,1].$ (5.16)
The next lemma is a basic bound on the growth of $|\mathbf{u}|$ along $\gamma$, where $u$ solves (5.8): it can grow exponentially at most like the integral of $|\psi_{\gamma,\alpha}^{\prime}|$. This is classical and different formulations can be found for instance in [Sj19, §7.1] but we give a direct proof in our special case.
Lemma 10.
Assume that $u$ solves (5.8) and $\mathbf{u}$ is given by (5.9). Then for $\gamma$ satisfying (5.15) we have for $t\in[0,1]$,
$|\mathbf{u}(\gamma(t))|\leq C_{\gamma}\exp\Big{(}h^{-1}\int_{0}^{t}|\psi^{\prime}_{\gamma,\alpha}(s)|\,ds\Big{)}|\mathbf{u}(\gamma(0))|.$ (5.17)
Proof.
For $u$ which solves (5.8), define the energy
$E(z):=\tfrac{1}{2}\big{(}|h\partial_{z}u(z)|^{2}+|\alpha W(z)u(z)|^{2}\big{)},\quad E_{\gamma}(t):=E(\gamma(t)).$ (5.18)
We may assume that $u$ is not identically 0 so that $E_{\gamma}(t)>0$ for all $t$. We compute
$\partial_{t}\log E_{\gamma}(t)$ $=2E_{\gamma}(t)^{-1}\operatorname{Re}\big{(}\partial_{z}E(\gamma(t))\gamma^{\prime}(t)\big{)},$
$\partial_{z}E(z)$ $=ih^{-1}\alpha W(z)\operatorname{Im}\big{(}\overline{\alpha W(z)u(z)}\cdot h\partial_{z}u(z)\big{)}+\tfrac{1}{2}\overline{W}(z)\partial_{z}W(z)|\alpha u(z)|^{2}.$
For all $z=\gamma(t)$, $0\leq t\leq 1$, we have
$\big{|}\operatorname{Im}\big{(}\overline{\alpha W(z)u(z)}\cdot h\partial_{z}u(z)\big{)}\big{|}$ $\leq E_{\gamma}(z),$
$\big{|}\tfrac{1}{2}\overline{W}(z)\partial_{z}W(z)|\alpha u(z)|^{2}\big{|}$ $\leq C_{\gamma}E_{\gamma}(z).$
This implies
$|\partial_{t}\log E_{\gamma}(t)|\leq 2h^{-1}\big{|}\operatorname{Im}\big{(}\alpha W(\gamma(t))\gamma^{\prime}(t)\big{)}\big{|}+C_{\gamma}.$
Integrating and using the identity
$\psi^{\prime}_{\gamma,\alpha}(t)=\operatorname{Im}\big{(}\alpha W(\gamma(t))\gamma^{\prime}(t)\big{)}$
we obtain
$|\log E_{\gamma}(t)-\log E_{\gamma}(0)|\leq C_{\gamma}+2h^{-1}\int_{0}^{t}|\psi^{\prime}_{\gamma,\alpha}(s)|\,ds\quad\text{for all }t\in[0,1].$
Since $C_{\gamma}^{-1}|\mathbf{u}(\gamma(t))|^{2}\leq E_{\gamma}(t)\leq C_{\gamma}|\mathbf{u}(\gamma(t))|^{2}$, this shows (5.17). ∎
5.2.2. WKB solutions
We next construct approximate WKB solutions, $u^{\pm}_{N}$, to the equation (5.8). Let $\gamma$ be a curve satisfying (5.15). Since $\gamma$ does not self ntersect, we can fix a simply connected open set set $\mathcal{U}_{\gamma}$ such that
$\gamma\ \subset\ \mathcal{U}_{\gamma}\ \subset\ \mathbb{C}\setminus W^{-1}(0).$
Our solutions have the form (where $N\geq 0$ is an integer)
$\begin{gathered}u^{\pm}_{N}(z)=\exp(i\Phi^{\pm}_{N}(z;\alpha,h)/h),\quad z\in\mathcal{U}_{\gamma},\\ \Phi^{\pm}_{N}(z;\alpha,h)=\pm\alpha\Phi(z)-ihb^{\pm}_{N}(z;h/\alpha),\end{gathered}$
where $\Phi$ is defined in (5.14). This means that
$u^{\pm}_{N}(z)=e^{\pm i\alpha\Phi(z)/h}A_{\pm}(z,h/\alpha),\ \ \ A_{\pm}(z,\tilde{h})=e^{b^{\pm}_{N}(z;\tilde{h})},\ \ \tilde{h}:=h/\alpha.$
The function whose exponential gives the amplitude of $u_{N}^{\pm}$ is given by
$b^{\pm}_{N}(z;\tilde{h})=\sum_{j=0}^{N}\tilde{h}^{j}a^{\pm}_{j}(z),$ (5.19)
where the coefficients $a^{\pm}_{j}\in\mathscr{O}(\mathcal{U}_{\gamma})$ are determined by the equation for $b^{\pm}_{N}$:
$2W(z)\partial_{z}b^{\pm}_{N}(z;\tilde{h})=-\partial_{z}W(z)\pm i\tilde{h}\big{(}\partial_{z}^{2}b^{\pm}_{N}(z;\tilde{h})+(\partial_{z}b^{\pm}_{N}(z;\tilde{h}))^{2}\big{)}+\mathcal{O}(\tilde{h}^{N+1}).$ (5.20)
Note that (5.20) determines the derivatives $\partial_{z}a^{\pm}_{j}(z)$ uniquely. Since $\partial_{z}a_{0}^{\pm}=-\frac{1}{2}\partial_{z}W/W$, we can take
$a^{\pm}_{0}(z)=-\tfrac{1}{2}\log W(z)$ (5.21)
where we fix a branch of $\log W$ on $\mathcal{U}_{\gamma}$.
Lemma 11.
In the notation of (5.16), we have for all $t\in[0,1]$ and $0<h<C_{\gamma,N}^{-1}$
$C_{\gamma,N}^{-1}e^{\mp\psi_{\gamma,\alpha}(t)/h}\leq|u^{\pm}_{N}(\gamma(t))|\leq C_{\gamma,N}e^{\mp\psi_{\gamma,\alpha}(t)/h},$ (5.22)
$|P(\alpha,h)u^{\pm}_{N}(\gamma(t))|\leq C_{\gamma,N}e^{\mp\psi_{\gamma,\alpha}(t)/h}h^{N+2}.$ (5.23)
Here $C_{\gamma,N}$ denotes a constant depending only on $N$, $\gamma$, $W$, and the constant $C_{0}$ in (5.7).
Proof.
The bound (5.22) follows immediately from the definitions of $u^{\pm}_{N}$ and $\psi_{\gamma,\alpha}$. To see (5.23), we compute
$P(\alpha,h)u^{\pm}_{N}(z)=$ $u^{\pm}_{N}(z)\big{(}(\partial_{z}\Phi^{\pm}_{N}(z;\alpha,h))^{2}-ih\partial_{z}^{2}\Phi^{\pm}_{N}(z;\alpha,h)-\alpha^{2}W(z)^{2}\big{)}$
$=$ $u^{\pm}_{N}(z)\big{(}\mp i\alpha h(2W(z)\partial_{z}b^{\pm}_{N}(z;h/\alpha)+\partial_{z}W(z))$
$-h^{2}(\partial_{z}^{2}b^{\pm}_{N}(z;h/\alpha)+(\partial_{z}b^{\pm}_{N}(z;h/\alpha))^{2})\big{)}.$
Together with (5.20) this gives (5.23). ∎
We also have the bound
$\partial_{z}\big{(}e^{b^{+}_{N}(z;\tilde{h})+b^{-}_{N}(z;\tilde{h})}\big{(}2W(z)-i\tilde{h}\partial_{z}(b^{+}_{N}(z;\tilde{h})-b^{-}_{N}(z;\tilde{h}))\big{)}\big{)}=\mathcal{O}(\tilde{h}^{N+1}),$ (5.24)
where $\tilde{h}=h/\alpha$. In fact, we have,
$\begin{split}\operatorname{Wr}(u_{N}^{+},u_{N}^{-})&=e^{b_{N}^{+}+b_{N}^{-}}\left(h\partial_{z}(\alpha\Phi-ihb_{N}^{+})-(-h\partial_{z}(\alpha\Phi-ihb_{N}^{-})\right)\\ &=\alpha e^{b_{N}^{+}+b_{N}^{-}}(2W-i\tilde{h}\partial_{z}(b_{N}^{+}-b_{N}^{-})).\end{split}$
On the other hand, (5.12), (5.22) and (5.23) give $\operatorname{Wr}(u_{N}^{+},u_{N}^{-})=\mathcal{O}(\tilde{h}^{N+1})$.
Next, using the WKB solutions, for a solution $u$ to (5.8) we define the vector of modified Wronskians
${\rm{\bf{Wr}}}_{N}(z)=\begin{pmatrix}u^{+}_{N}(z)^{-1}\operatorname{Wr}(u,u^{+}_{N})(z)\\ u^{-}_{N}(z)^{-1}\operatorname{Wr}(u,u^{-}_{N})(z)\end{pmatrix},\quad z\in\mathcal{U}_{\gamma}.$ (5.25)
Then, for all $t\in[0,1]$,
${\rm{\bf{Wr}}}_{N}(z)=B_{N}(z;\alpha,h)\mathbf{u}(z)$ (5.26)
where
$B_{N}(z;\alpha,h)=\begin{pmatrix}\partial_{z}\Phi^{+}_{N}(z;\alpha,h)&-1\\ \partial_{z}\Phi^{-}_{N}(z;\alpha,h)&-1\end{pmatrix}=\begin{pmatrix}\ \ \alpha W(z)+\mathcal{O}(h)&-1\\ -\alpha W(z)+\mathcal{O}(h)&-1\end{pmatrix}$. (5.27)
Since $W\neq 0$, both $B_{N}$ and its inverse are bounded in norm uniformly in $h$ for small $h$.
5.2.3. Approximating the transition matrix
We now assume that $\gamma$ is a curve as in (5.15) and $\gamma,\alpha$ satisfy the additional condition
$\psi_{\gamma,\alpha}^{\prime}(t)\geq 0\quad\text{for all }t\in[0,1],$ (5.28)
where the function $\psi_{\gamma,\alpha}$ was defined in (5.16). Define the diagonal matrix
$S_{\gamma}(\alpha,h):=\begin{pmatrix}e^{i(\Phi_{+}(\gamma(0);\alpha,h)-\Phi_{+}(\gamma(1);\alpha,h))/h}&0\\ 0&e^{i(\Phi_{-}(\gamma(0);\alpha,h)-\Phi_{-}(\gamma(1);\alpha,h))/h}\end{pmatrix}$. (5.29)
The transition matrix $\mathcal{M}_{\gamma(1)\leftarrow\gamma(0)}$ defined in (5.10) is now approximated as follows:
Lemma 12.
Under the condition (5.28), and with $B_{N}$ defined in (5.27), we have
$\mathcal{M}_{\gamma(1)\leftarrow\gamma(0)}=$ $\,B_{N}(\gamma(1);\alpha,h)^{-1}S_{\gamma}(\alpha,h)B_{N}(\gamma(0);\alpha,h)$ (5.30)
$+\mathcal{O}_{\gamma,N}(h^{N+1}e^{(\psi_{\gamma,\alpha}(1)-\psi_{\gamma,\alpha}(0))/h}).$
Proof.
1. Let $u$ be a solution to (5.8). By Lemma 10 and (5.28), we have for all $t\in[0,1]$
$|\mathbf{u}(\gamma(t))|\leq C_{\gamma}e^{(\psi_{\gamma,\alpha}(t)-\psi_{\gamma,\alpha}(0))/h}|\mathbf{u}(\gamma(0))|.$ (5.31)
Define the Wronskians
$\operatorname{Wr}_{\pm}(t):=\operatorname{Wr}(u,u^{\pm}_{N})(\gamma(t)),\quad{\rm{\bf{Wr}}}_{N}(\gamma(t)):=\begin{pmatrix}e^{-i\Phi^{+}_{N}(\gamma(t);\alpha,h)/h}\operatorname{Wr}_{+}(t)\\ e^{-i\Phi^{-}_{N}(\gamma(t);\alpha,h)/h}\operatorname{Wr}_{-}(t)\end{pmatrix}$.
By (5.12) and since $P(\alpha,h)u=0$, we have for all $t\in[0,1]$
$|\partial_{t}\operatorname{Wr}_{\pm}(t)|$ $\leq C_{\gamma}h^{-1}|u(\gamma(t))|\cdot|P(\alpha,h)u^{\pm}_{N}(\gamma(t)|$
$\leq C_{\gamma,N}h^{N+1}e^{(\mp\psi_{\gamma,\alpha}(t)+\psi_{\gamma,\alpha}(t)-\psi_{\gamma,\alpha}(0))/h}|\mathbf{u}(\gamma(0))|$
where the last inequality follows from (5.23) and (5.31). This implies that
$|\partial_{t}\operatorname{Wr}_{+}(t)|$ $\leq C_{\gamma,N}h^{N+1}e^{-\psi_{\gamma,\alpha}(0)/h}|\mathbf{u}(\gamma(0))|,$
$|\partial_{t}\operatorname{Wr}_{-}(t)|$ $\leq C_{\gamma,N}h^{N+1}e^{(2\psi_{\gamma,\alpha}(1)-\psi_{\gamma,\alpha}(0))/h}|\mathbf{u}(\gamma(0))|.$
Integrating in $t\in[0,1]$, we see that
$\operatorname{Wr}_{+}(1)$ $=\operatorname{Wr}_{+}(0)+\mathcal{O}_{\gamma,N}(h^{N+1}e^{-\psi_{\gamma,\alpha}(0)/h}|\mathbf{u}(\gamma(0))|),$ (5.32)
$\operatorname{Wr}_{-}(1)$ $=\operatorname{Wr}_{-}(0)+\mathcal{O}_{\gamma,N}(h^{N+1}e^{(2\psi_{\gamma,\alpha}(1)-\psi_{\gamma,\alpha}(0))/h}|\mathbf{u}(\gamma(0))|).$
2. Multiplying both sides of (5.32) by $e^{-i\Phi^{\pm}_{N}(\gamma(1);\alpha,h)/h}$, we get
$e^{-i\Phi_{+}(\gamma(1);\alpha,h)/h}\operatorname{Wr}_{+}(1)$ $=e^{(-i\Phi_{+}(\gamma(1);\alpha,h)+i\Phi_{+}(\gamma(0);\alpha,h))/h)}e^{-i\Phi_{+}(\gamma(0);\alpha,h)}\operatorname{Wr}_{+}(t)$
$\ \ \ \ \ \ +\mathcal{O}_{\gamma,N}(h^{N+1}e^{\psi_{\gamma,\alpha}(1)/h-\psi_{\gamma,\alpha}(0)/h}|\mathbf{u}(\gamma(0))|),$
$e^{-i\Phi_{-}(\gamma(1);\alpha,h)/h}\operatorname{Wr}_{-}(1)$ $=e^{(-i\Phi_{-}(\gamma(1);\alpha,h)+i\Phi_{-}(\gamma(0);\alpha,h))/h)}e^{-i\Phi_{-}(\gamma(0);\alpha,h)}\operatorname{Wr}_{+}(t)$
$\ \ \ \ \ \ +\mathcal{O}_{\gamma,N}(h^{N+1}e^{\psi_{\gamma,\alpha}(1)/h-\psi_{\gamma,\alpha}(0)/h}|\mathbf{u}(\gamma(0))|),$
(we used here the facts that $\operatorname{Re}(-i\Phi_{\pm}(\gamma(1);\alpha,h)=\pm\psi_{\gamma,\alpha}(1)+\mathcal{O}(h)$) that is,
${\rm{\bf{Wr}}}_{N}(\gamma(1))=S_{\gamma}(\alpha,h){\rm{\bf{Wr}}}_{N}(\gamma(0))+\mathcal{O}_{\gamma,N}(h^{N+1}e^{(\psi_{\gamma,\alpha}(1)-\psi_{\gamma,\alpha}(0))/h}|\mathbf{u}(\gamma(0))|).$
By (5.26) this implies
$B_{N}(\gamma(1);\alpha,h)\mathbf{u}(\gamma(1))=$ $S_{\gamma}(\alpha,h)B_{N}(\gamma(0);\alpha,h)\mathbf{u}(\gamma(0))$
$+\mathcal{O}_{\gamma,N}(h^{N+1}e^{(\psi_{\gamma,\alpha}(1)-\psi_{\gamma,\alpha}(0))/h}|\mathbf{u}(\gamma(0))|).$
Therefore
$\mathbf{u}(\gamma(1))=$ $B_{N}(\gamma(1);\alpha,h)^{-1}S_{\gamma}(\alpha,h)B_{N}(\gamma(0);\alpha,h)\mathbf{u}(\gamma(0))$
$+\mathcal{O}_{\gamma,N}(h^{N+1}e^{(\psi_{\gamma,\alpha}(1)-\psi_{\gamma,\alpha}(0))/h}|\mathbf{u}(\gamma(0))|).$
Since this holds for each solution $u$ to (5.8), we get (5.30). ∎
5.3. WKB analysis in the periodic case
From now on, we assume that the potential $W$ is 1-periodic:
$W(z+1)=W(z).$
We also assume that $\gamma$ is a curve as in (5.15) such that
$\gamma(1)=\gamma(0)+1,$ (5.33)
and $W$ has winding number 0 along $\gamma$:
$\int_{\gamma}{W^{\prime}(z)\over W(z)}\,dz=0.$ (5.34)
5.3.1. Periodicity of WKB solutions
Let $u^{\pm}_{N}(z)=\exp(i\Phi^{\pm}_{N}(z;\alpha,h)/h)$, $z\in\mathcal{U}_{\gamma}$ (a simply connected neighbourhood of $\gamma$ which satisfies (5.15) and (5.33)) be the approximate WKB solutions to (5.8) constructed in §5.2.2. Shrinking $\mathcal{U}_{\gamma}$ if necessary, we assume that the set $\mathcal{U}_{\gamma}\cap(\mathcal{U}_{\gamma}-1)$ is connected.
The solutions $u^{\pm}_{N}$ have the following periodicity property:
Lemma 14.
There exist constants $c_{0},\dots,c_{N+1}$ such that for all $z\in\mathcal{U}_{\gamma}\cap(\mathcal{U}_{\gamma}-1)$
$\begin{gathered}\Phi^{\pm}_{N}(z+1;\alpha,h)-\Phi^{\pm}_{N}(z;\alpha,h)=\pm\alpha c(h/\alpha),\\ c(\tilde{h})=\sum_{j=0}^{N+1}c_{j}\tilde{h}^{j},\quad c_{0}=\int_{0}^{1}W(z)\,dz,\quad c_{1}=0.\end{gathered}$ (5.35)
Proof.
Since $\partial_{z}\Phi=W$ and $W$ is 1-periodic, we have $\partial_{z}(\Phi(z+1)-\Phi(z))=0$. Therefore, $\Phi(z+1)-\Phi(z)$ is constant:
$\Phi(z+1)-\Phi(z)=\int_{0}^{1}W(z)dz=:c_{0}.$
[…]
$\partial_{z}a^{\pm}_{j}(z+1)=\partial_{z}a^{\pm}_{j}(z),\quad z\in\mathcal{U}_{\gamma}\cap(\mathcal{U}_{\gamma}-1).$ (5.36)
It follows that
$a^{\pm}_{j}(z+1)-a^{\pm}_{j}(z)=c^{\pm}_{j},\quad z\in\mathcal{U}_{\gamma}\cap(\mathcal{U}_{\gamma}-1)$ (5.37)
for some constants $c^{\pm}_{j}$.
Next, by (5.24) we have for all $z\in\mathcal{U}_{\gamma}\cap(\mathcal{U}_{\gamma}-1)$
$\begin{gathered}e^{b^{+}_{N}(z+1;\tilde{h})+b^{-}_{N}(z+1;\tilde{h})}\big{(}2W(z+1)-i\tilde{h}\partial_{z}(b^{+}_{N}(z+1;\tilde{h})-b^{-}_{N}(z+1;\tilde{h}))\big{)}\\ =e^{b^{+}_{N}(z;\tilde{h})+b^{-}_{N}(z;\tilde{h})}\big{(}2W(z)-i\tilde{h}\partial_{z}(b^{+}_{N}(z;\tilde{h})-b^{-}_{N}(z;\tilde{h}))\big{)}+\mathcal{O}(\tilde{h}^{N+1}).\end{gathered}$
By (5.36) the terms in the parentheses on both sides are the same, thus
$e^{b^{+}_{N}(z+1;\tilde{h})+b^{-}_{N}(z+1;\tilde{h})}=e^{b^{+}_{N}(z;\tilde{h})+b^{-}_{N}(z;\tilde{h})}+\mathcal{O}(\tilde{h}^{N+1}).$
In other words,
$\exp\Big{(}\sum_{j=0}^{N}\tilde{h}^{j}(c_{j}^{+}+c_{j}^{-})\Big{)}=1+\mathcal{O}(\tilde{h}^{N+1}).$ (5.38)
By (5.21) $a^{\pm}_{0}(z)=\frac{1}{2}\log W(z)$ and hence $c_{0}^{\pm}=\frac{1}{2}\int_{\gamma}W^{\prime}/Wdz$. The winding condition (5.34) then shows that $c^{+}_{0}+c^{-}_{0}=0$. Expanding the exponential in (5.38) then gives
$c^{+}_{j}+c^{-}_{j}=0.$
Recalling that $\Phi_{N}^{\pm}=\pm\alpha\Phi-i\sum_{j=0}^{N}\tilde{h}^{j}a_{j}^{\pm}$, gives (5.35) with $c_{j}:=-ic^{+}_{j-1}$, $j\geq 1$. In view of (5.21), $c_{1}=\frac{i}{2}\int_{\gamma}W^{\prime}(z)/W(z)dz=0$, by assumption (5.34). ∎
5.3.2. Stokes loops
We now introduce the notion of a Stokes loop. If the potential $W$ admits a Stokes loop then we are able to prove a quantization condition for the corresponding operator $P$.
Definition 15.
Assume that $W$ is a 1-periodic entire function on $\mathbb{C}$. Let $\gamma:[0,1]\to\mathbb{C}$ be a smooth non-self-intersecting curve. We say that $\gamma$ is a Stokes loop for $W$, if $\gamma$ extends to a smooth non-self-intersecting curve $\widetilde{\gamma}$$\mathbb{{\mathbb{R}}}\to\mathbb{C}$ and the following conditions hold:
$\widetilde{\gamma}(t+1)$ $=\widetilde{\gamma}(t)+1,$ (5.39)
$W(\gamma(t))\gamma^{\prime}(t)$ $>0\quad\text{for all }t\in[0,1]$ (5.40)
Another way to formulate the condition on the curve $\gamma$ is to consider it as a closed non-self-intersecting curve $\gamma:\mathbb{R}/\mathbb{Z}\mapsto\mathbb{C}/\mathbb{Z}$, where the action is $z\mapsto z+1$. Indeed, $\widetilde{\gamma}$ defines such a curve. That curve is non-self-intersecting: $\gamma([s])=\gamma([t])$ ($t\mapsto[t]$ is the map $\mathbb{R}\to\mathbb{R}/\mathbb{Z}$) means that for some $m$, $\widetilde{\gamma}(s)=\widetilde{\gamma}(t)+m=\widetilde{\gamma}(t+m)$. As $\widetilde{\gamma}$ is non-self-intersecting, this implies that $s=t+m$, that is $[s]=[t]$.
It is clear that a Stokes loop satisfies (5.15). We also have (5.34):
Lemma 16.
Suppose that $\gamma$ is a Stokes loop in the sense of Definition 15. Then the the winding number condition (5.34) holds.
Proof.
It is convenient to change the parametrisation of $\gamma$ to $\rho:[0,L]\to\mathbb{C}$, $|\rho^{\prime}(s)|=1$, so that now $\gamma(L)=\gamma(0)+1$. Similarly it is useful to define $w(s):=W(\rho(s))/|W(\rho(s))|$ and to note that the winding numbers around 0 of $s\mapsto W(\rho(s))$ and of $s\mapsto w(s)$ are the same. The condition (5.40) becomes $w(s)\rho^{\prime}(s)=1$, that is
$w(s)=\overline{\rho^{\prime}(s)}.$ (5.41)
Condition (5.39) shows that $s\mapsto\rho^{\prime}(s)$ is a smooth closed curve. Also, (5.41) shows that the winding number around zero of the curve $w$ is the negative of the same number for the curve $\rho^{\prime}$.
Hence it suffices to show that the winding number of $0,L]\ni s\mapsto\rho^{\prime}(s)$ is 0. As remarked after Definition 15 we can consider $\rho$ as a closed curve on $\mathbb{C}/\mathbb{Z}$. Let $0,L]\ni s\mapsto\theta(s)$ be continuously chosen angle between $\rho^{\prime}(s)$ and 11 (thought of as vectors in $\mathbb{R}^{2}\simeq\mathbb{C}$). The curvature of $\rho$ is given by $\kappa_{\rho}(s)=\theta^{\prime}(s)$. (We recall that to see this we write $\rho^{\prime}(s)=(\cos\theta(s),\sin\theta(s))$ so that $\kappa_{\rho}(s)=\rho^{\prime\prime}(s)\cdot\rho^{\prime}(s)^{\perp}=\theta^{\prime}(s)$). Let $\rho_{1}:[0,1]\to\mathbb{C}/\mathbb{Z}$ be given by $\rho_{1}(s)=-s+iM\mod 1$ with $M$ chosen large enough so that $\rho_{1}$ and $\rho$ are disjoint. Then the curves $\rho_{1}$ and $\rho$ form a boundary of a bounded subset of $\mathbb{C}/\mathbb{Z}$ homeomorphic to $(\mathbb{R}/\mathbb{Z})\times[0,1]$ (this follows from the Jordan–Schönflies theorem – see [NaRa12, Theorem 6.7.1] – applied to the cylinder compactified at one of the infinities) and the Gauss–Bonnet theorem for a flat surface (the cylinder $\mathbb{C}/\mathbb{Z}$; see for instance [DoCa16, §4-5] for the general case) shows that
$\int_{\rho}\kappa_{\rho}(s)ds+\int_{\rho_{1}}\kappa_{\rho_{1}}(s)ds=0.$
Since $\kappa_{\rho_{1}}\equiv 0$, it follows that $\int_{\rho}\kappa_{\rho}(s)ds=0$. The calculation of $\kappa_{\rho}(s)$ above also shows that $\kappa_{\rho}(s)=-i\overline{\rho^{\prime}(s)}\rho^{\prime\prime}(s)=-i\rho^{\prime\prime}(s)/\rho^{\prime}(s)$. The winding number of $0,L]\ni s\mapsto\rho^{\prime}(s)$ is
$\frac{1}{2\pi i}\int_{0}^{L}\rho^{\prime\prime}(s)/\rho^{\prime}(s)ds=\frac{1}{2\pi}\int_{\rho}\kappa_{\rho}(s)ds=0,$
which proves the claim. ∎
Moreover, any Stokes loop $\gamma$ satisfies the condition (5.28) for $\alpha$ in the upper half-plane:
$\psi_{\gamma,\alpha}^{\prime}(t)=\operatorname{Im}\big{(}\alpha W(\gamma(t))\gamma^{\prime}(t)\big{)}\geq 0\quad\text{for all }t\in[0,1],\ \operatorname{Im}\alpha\geq 0.$ (5.42)
We also get that the coefficient $c_{0}$ defined in (5.35) is positive:
$c_{0}:=\int_{\gamma}W(z)\,dz=\int_{0}^{1}W(\gamma(t))\gamma^{\prime}(t)\,dt>0.$ (5.43)
If $W=c_{0}>0$ is a constant potential, then it has a Stokes loop $\gamma(t)=t$. The next lemma shows that perturbations of this potential have Stokes loops as well:
Proposition 17.
Assume that $W$ is a 1-periodic entire function and
$\int_{0}^{1}W(z)\,dz=c_{0}>0,$ (5.44)
$W(z)\neq 0\quad\text{and}\quad|\operatorname{Im}W(z)|\leq\operatorname{Re}W(z)\quad\text{when }|\operatorname{Im}z|\leq 1.$ (5.45)
Then $W$ has a Stokes loop.
Proof.
We are looking for $\gamma(t)$ of the form
$\gamma(t):=t+if(t),\ \ f\in C^{\infty}(\mathbb{R};\mathbb{R}).$ (5.46)
The condition (5.40) then means that
$\begin{split}&\operatorname{Re}W(t+if(t))-f^{\prime}(t)\operatorname{Im}W(t+if(t))>0,\\ &f^{\prime}(t)\operatorname{Re}W(t+if(t))+\operatorname{Im}W(t+if(t))=0.\end{split}$
This means that $f$ has to solve
$f^{\prime}(t)=-{\operatorname{Im}W(t+if(t))\over\operatorname{Re}W(t+if(t))},$
and we choose $f(0)=0$ as the initial condition. By (5.45) this Cauchy problem has a solution for $0\leq t\leq 1$ and and moreover $|f(t)|\leq 1$ there. In the notation of (5.46) we have
$W(\gamma(t))\gamma^{\prime}(t)={|W(t+if(t))|^{2}\over\operatorname{Re}W(t+if(t))}>0,\ \ \ t\in[0,1].$
In particular,
$\int_{0}^{1+if(1)}W(z)\,dz=\int_{0}^{1}W(\gamma(t))\gamma^{\prime}(t)\,dt>0.$
Since $\int_{0}^{1}W(z)\,dz>0$, we see that
$0=\operatorname{Im}\int_{1}^{1+if(1)}W(z)\,dz=\int_{0}^{f(1)}\operatorname{Re}W(1+is)\,ds.$
Since $\operatorname{Re}W(1+is)>0$, we see that $f(1)=0$. But then periodicity of $W$ shows that $f_{1}(t):=f(t+1)$ solves the same equation as $f(t)$ with the same initial data. Hence $f(t)=f(t+1)$ and $\gamma(t+1)=\gamma(t)+1$. Thus $\gamma$ is a Stokes loop. ∎
5.4. Mutliplicity in terms of the transition matrix
For $k\in\mathbb{C}$, let $\mathcal{V}_{k,h}$ be the space of entire functions satisfying the periodicity condition
$u(z+1)=e^{ik/h}u(z),\quad z\in\mathbb{C}.$ (5.47)
We study the values of $(\alpha,k)$ for which the operator $P(\alpha,h)$ has a kernel on $\mathcal{V}_{k,h}$. More precisely, for $\alpha$ satisfying (5.7) and a fixed $k\in\mathbb{C}$, we define the multiplicity in the way similar to (2.7) (but with the roles of $\alpha$ and $k$ reversed). For that we consider
$\begin{gathered}P_{k}(\alpha,h):H^{2}_{k}\to L^{2}_{k},\ \ \ H^{s}_{k}:=\{w\in L^{s}_{\rm{loc}}(\mathbb{R}):w(x+1)=e^{ik/h}w(x)\},\end{gathered}$ (5.48)
$(/H_{k}^{0}=L^{2}_{k}=L^{2}$($\mathbb{R}$$\mathbb{Z}$) and then put
$m_{h}(\alpha,k)=\frac{1}{2\pi i}\operatorname{tr}_{L^{2}_{k}}\oint_{\alpha}P_{k}(\zeta,h)^{-1}\partial_{\zeta}P_{k}(\zeta,h)d\zeta,$ (5.49)
where the integral is over a positively oriented circle including $\alpha$ as the only possible pole of $\zeta\mapsto P_{k}(\zeta,h)$. Since, in the notation of (1.1), $P_{k}(\zeta,h)=h^{2}P(\zeta/h,k)$ we see that
$m_{h}(\alpha,k)=m(\alpha/h,k/h),$ (5.50)
where the right hand side was defined in (5.6).
We note that, in the notation of (1.13)
$\begin{gathered}P_{k}(\alpha,h)=h^{2}\tau(k)P_{k/h}(\alpha/h)\tau(k)^{-1},\\ \tau(k)w(x):=e^{ikx/h}w(x),\ \ \ \tau(k):H^{s}(\mathbb{R}/\mathbb{Z})\to H^{s}_{k}.\end{gathered}$
Hence, the meromorphy of $\zeta\mapsto P_{k}(\zeta,h)$ is guaranteed by Lemma 9, except for $k/h\in 2\pi\mathbb{Z}$ when we can have $\ker P_{k}(\alpha,h)\neq\{0\}$ for all $\alpha$ (see the remark after Lemma 9). In that case, we put $m_{h}(\alpha,k)=\infty$.
We will now relate the multiplicity defined by (5.49) to the multiplicity of zeros of an eigenvalue equation for $\mathcal{M}_{z_{0}+1\leftarrow z_{0}}$. We note that for fixed $h$, the holomorphic dependence on the parameter $\alpha$ shows that
$\mathbb{C}\ni\alpha\mapsto\mathcal{M}_{z_{0}+1\leftarrow z_{0}}(\alpha,h)\ \text{ is holomorphic.}$
Proposition 18.
Fix $k\in\mathbb{C}$ and $h>0$. For arbitrary $z_{0}\in\mathbb{C}$
$\alpha\mapsto P_{k}(\alpha,h)^{-1}\text{ is meromorphic }\Longleftrightarrow\ \det(\mathcal{M}_{z_{0}+1\leftarrow z_{0}}(\alpha,h)-e^{ik})\not\equiv 0.$
If the determinant on the right hand side is not identically 0, then
$m_{h}(\alpha,k)=\frac{1}{2\pi i}\operatorname{tr}_{\mathbb{C}^{2}}\oint_{\alpha}(\mathcal{M}_{z_{0}+1\leftarrow z_{0}}(\zeta,h)-e^{ik/h})^{-1}\partial_{\zeta}\mathcal{M}_{z_{0}+1\leftarrow z_{0}}(\zeta,h)d\zeta,$ (5.51)
where the integral is as in (5.49). In other words, the multiplicity $m_{h}(\alpha,k)$ is the same as the multiplicity of the zero of the eigenvalue equation
$\zeta\mapsto\det(\mathcal{M}_{z_{0}+1\leftarrow z_{0}}(\zeta,h)-e^{ik/h})$
at $\zeta=\alpha$.
Proof.
We start by observing that since $W(z+1)=W(z)$, $\mathcal{M}_{z_{0}\leftarrow z}=\mathcal{M}_{z_{0}+1\leftarrow z+1}$. Hence
$\mathcal{M}_{z+1\leftarrow z}=\mathcal{M}_{z+1\leftarrow z_{0}+1}\mathcal{M}_{z_{0}+1\leftarrow z_{0}}\mathcal{M}_{z_{0}\leftarrow z}=\mathcal{M}_{z\leftarrow z_{0}}\mathcal{M}_{z_{0}+1\leftarrow z_{0}}\mathcal{M}_{z\leftarrow z_{0}}^{-1}$ (5.52)
and the statements for a fixed $z_{0}$ are equivalent to statements for all $z$. The obstruction to the meromorphy of $P_{k}(\alpha,h)^{-1}$ is having a non-trivial solution $w(\alpha)\in H_{k}^{2}$ to $P_{k}(\alpha,h)w(\alpha)=0$ for all $\alpha$. As discussed in §5.1 $x\mapsto w(\alpha,x)$ extends to an entire function which then satisfies (5.47) and $P(\alpha,h)w=0$. Definition of $\mathcal{M}_{z+1\leftarrow z}(\alpha,h)$ in §5.1, then shows that $e^{ik/h}$ is an eigenvalue of $\mathcal{M}_{z+1\leftarrow z}$ for all $z$. That gives $\Leftarrow$ in the first statement. To see $\Rightarrow$ we note that meromorphy of $P_{k}(\alpha,h)^{-1}$ implies existence of $\alpha_{0}$ for which $\ker P_{k}(\alpha_{0},h)=\{0\}$. If $e^{ik/h}$ were an eigenvalue of $\mathcal{M}_{z_{0}+1\leftarrow z_{0}}(\alpha_{0},h)$ then, by (5.52) it would be an eigenvalue of $\mathcal{M}_{z+1\leftarrow z}(\alpha_{0},h)$ for all $z$ and we would have have an entire solution to $P(\alpha_{0},h)u=0$ satisfying (5.47). But that means that $\ker P_{k}(\alpha,h)$ is nontrivial.
The formula (5.51) makes the above argument quantitative and it is easier to obtain in the setting of operators acting on periodic function:
$\widetilde{P}_{k}(\alpha,h):=e^{-ixk/h}P_{k}(\alpha,h)e^{ixk/h}:H^{2}(\mathbb{R}/\mathbb{Z})\to L^{2}(\mathbb{R}/\mathbb{Z}).$
The related operator acting on holomorphic functions is given by $\widetilde{P}_{k}(\alpha,h)=(hD_{z}+k)^{2}-\alpha^{2}$. The corresponding transition matrix, $\mathcal{M}^{k}_{z_{1}\leftarrow z_{0}}$, is defined as in (5.10) but with (5.9) replaced by $\mathbf{u}(z)=(u(z),(hD_{z}+k)u(z))$. It is related to the original transition matrix as follows:
$\mathcal{M}^{k}_{z_{1}\leftarrow z_{0}}(\alpha,h)=e^{-ik(z_{0}-z_{1})/h}\mathcal{M}_{z_{1}\leftarrow z_{0}}(\alpha,h).$ (5.53)
The final reduction is to a system:
$D_{k}(\alpha,h):=\begin{pmatrix}hD_{z}+k&-1\\ -V&hD_{z}+k\end{pmatrix},\ \ \ D_{k}(\alpha,h)\mathcal{M}^{k}_{z\leftarrow z_{0}}(\alpha,h)=0,$ (5.54)
and $\mathcal{M}^{k}_{z_{0}\leftarrow z_{0}}(\alpha,h)=I_{\mathbb{C}^{2}}$.
Proceeding as in (2.13) we see that in the definition of $m_{h}(\alpha,k)$ we can replace $P_{k}(\alpha,h)$ by $D_{k}(\alpha,h)$ (now acting on periodic $\mathbb{C}^{2}$-valued functions) and in view of (5.52) and (5.53), the equality (5.51) is equivalent to
$\begin{split}&\frac{1}{2\pi i}\operatorname{tr}_{L^{2}(\mathbb{R}/\mathbb{Z};\mathbb{C}^{2})}\oint_{\alpha}D_{k}(\zeta,h)^{-1}\partial_{\zeta}D_{k}(\zeta,h)d\zeta\\ &\ \ \ \ \ \frac{1}{2\pi i}\operatorname{tr}_{\mathbb{C}^{2}}\oint_{\alpha}(\mathcal{M}^{k}_{1\leftarrow 0}(\zeta,h)-I)^{-1}\partial_{\zeta}\mathcal{M}^{k}_{1\leftarrow 0}(\zeta,h)d\zeta.\end{split}$ (5.55)
A systematic derivation of this formula follows for setting up a Grushin problem (see [DyZw19, §C.1]) for $D_{k}(\alpha,h)$ following [SZ07, (3.12)]. To pose it we introduce
$\chi\in C^{\infty}(\mathbb{R}/\mathbb{Z};[0,1]),\ \ \ \mathbb{R}/\mathbb{Z}\simeq(-\tfrac{1}{2},\tfrac{1}{2}],\ \ \ \operatorname{supp}\chi=[-\tfrac{1}{4},\tfrac{1}{2}],\ \ \ \chi|_{[-\frac{1}{8},\frac{1}{4}]}=1.$
We write $\chi^{\prime}=\chi^{\prime}_{+}+\chi_{-}^{\prime}$ where $\chi_{+}^{\prime}$ is supported near $\frac{1}{4}$ and $\chi_{-}^{\prime}$ is supported near $\pm\frac{1}{2}$. The following operators are also well defined:
$\begin{gathered}R_{-}:\mathbb{C}^{2}\to H^{2}(\mathbb{R}/\mathbb{Z}),\ \ \ (R_{-}u_{-})(x):=\chi^{\prime}_{-}\mathcal{M}^{k}_{x\leftarrow 0},\\ R_{+}:H^{2}(\mathbb{R}/\mathbb{Z})\to\mathbb{C}^{2},\ \ \ R_{+}u:=u(0).\end{gathered}$ (5.56)
We then claim that
$\mathscr{P}:=\begin{pmatrix}(i/h)D_{k}(\alpha,h)&R_{-}\\ R_{+}&0\end{pmatrix}:H^{1}(\mathbb{R}/\mathbb{Z};\mathbb{C}^{2})\oplus\mathbb{C}^{2}\to L^{2}(\mathbb{R}/\mathbb{Z};\mathbb{C}^{2})\oplus\mathbb{C}^{2},$ (5.57)
is invertible with the inverse given by
$\mathscr{P}^{-1}=\mathscr{E}:=\begin{pmatrix}E&E_{-}\\ E_{+}&E_{-+}\end{pmatrix},\ \ \ E_{-+}=I_{\mathbb{C}^{2}}-\mathcal{M}^{k}_{1\leftarrow 0}(\alpha,h).$ (5.58)
The conclusion (5.55), and hence (5.51), then follow from [SZ07, Proposition 4.1].
To show invertibility of (5.56) it is enough to construct a left inverse. The operator $D_{k}(\alpha,h)$ is a Fredholm operator and as
$\mathscr{P}^{t}_{k}(\alpha,h):=\begin{pmatrix}(i/h)D_{k}(\alpha,h)&(1-t)R_{-}\\ (1-t)R_{+}&tI_{\mathbb{C}^{2}}\end{pmatrix}:H^{1}(\mathbb{R}/\mathbb{Z};\mathbb{C}^{2})\oplus\mathbb{C}^{2}\to L^{2}(\mathbb{R}/\mathbb{Z};\mathbb{C}^{2})\oplus\mathbb{C}^{2},$
are finite rank perturbations of a Fredholm operator (with $D_{k}(\alpha,h)$ as the only non-zero entry), they are also Fredholm operators (see for instance [Hö88, §2.5] for basic properties of those operators). Since $D_{k}(0,k)$ is invertible for $k/h\notin 2\pi\mathbb{Z}$, it follows that $\mathscr{P}^{k}_{1}$ is invertible. Hence continuity of the index shows that the index of $\mathscr{P}$ is 0 and the right inverse is automatically the left inverse.
We construct the right inverse explicitly as follows. We first consider the problem
$(i/h)D_{k}\mathbf{u}(x)+\chi^{\prime}_{-}(x)\mathcal{M}^{k}_{x\leftarrow 0}u_{-}=0,\ \ \ \mathbf{u}(0)=v_{+}\in\mathbb{C}^{2}.$ (5.59)
We can define multivalued matrix valued functions on $\mathbb{R}/\mathbb{Z}$ as follows:
$I_{\pm}([x])=\mathcal{M}^{k}_{x\leftarrow 0},\ \ \pm x\geq-\tfrac{1}{4},$ (5.60)
where $x\mapsto[x]$ is the map $\mathbb{R}\to\mathbb{R}/\mathbb{Z}$. We can think of $I_{+}(x)$ as solving the equation forward from 0 and $I_{-}(x)$ backward from 0. The function $\chi(x)I_{+}(x)v_{+},(1-\chi(x))I_{-}(x)v_{+}$ are now well defined as functions on the circle (we drop $[x]$ notation when there is no ambiguity). We can then put
$\mathbf{u}(x)=\chi(x)I_{+}(x)v_{+}+(1-\chi(x))I_{-}(x)v_{+}\in H^{1}(\mathbb{R}/\mathbb{Z};\mathbb{C}^{2}),$
so that
$\begin{split}(i/h)D_{k}\mathbf{u}&=(i/h)[D_{k},\chi]I_{+}v_{+}-(i/h)[D_{k},\chi]I_{-}v=\chi^{\prime}I_{+}v_{+}-\chi^{\prime}I_{-}v_{+}\\ &=\chi^{\prime}_{-}(x)(I_{+}(x)v_{+}-I_{-}(x)v_{+}),\end{split}$
where we used the fact that $I_{+}(x)v_{+}=I_{-}(x)v_{+}$ on the support of $\chi_{+}^{\prime}$ (near 0). On the other hand for $[x]$ on the support of $\chi_{-}^{\prime}$, say for $x$ near $\frac{1}{2}]$, we have
$\begin{split}\chi^{\prime}_{-}([x])I_{+}([x])v_{+}&=\chi^{\prime}_{-}(x)\mathcal{M}^{k}_{x\leftarrow 0}=\chi_{-}^{\prime}(x)\mathcal{M}^{k}_{x-1\leftarrow-1}=\chi_{-}^{\prime}(x)\mathcal{M}^{k}_{x-1\leftarrow 0}\mathcal{M}^{k}_{0\leftarrow-1}\\ &=\chi_{-}^{\prime}(x)I_{-}([x])\mathcal{M}^{k}_{1\leftarrow 0}.\end{split}$
Hence (5.59) is solved by putting
$\mathbf{u}=E_{+}v_{+}:=(\chi I_{+}+(1-\chi)I_{-})v_{+},\ \ \ u_{-}=E_{-+}v_{+}:=(I_{\mathbb{C}^{2}}-\mathcal{M}^{k}_{1\leftarrow 0})v_{+}.$ (5.61)
It remains to solve
$(i/h)D_{k}\mathbf{u}(x)+\chi^{\prime}_{-}(x)\mathcal{M}^{k}_{x\leftarrow 0}u_{-}=v\in L^{2}(\mathbb{R}/\mathbb{Z};\mathbb{C}^{2}),\ \ \ \mathbf{u}(0)=0,$
and that is done similarly using the forward and backward solutions to the non-homogeneous problem. Since all that matters for us is the formula for $E_{-+}$ in (5.61), we leave the details to the reader. This completes the proof of (5.58) and as explained there of (5.51). ∎
5.5. Quantisation condition
Now, we assume that $\gamma$ is a Stokes loop and $\operatorname{Im}\alpha\geq 0$. We note that this assumption can be relaxed to $\operatorname{Im}\alpha\geq-h^{M}$ for some (large) $M$.
We will then put $z_{0}:=\gamma(0)$ in Proposition 18, so that $\mathcal{M}_{z_{0}+1\leftarrow z_{0}}=\mathcal{M}_{\gamma(1)\leftarrow\gamma(0)}$. Putting
$Z_{h}(\alpha):=\tfrac{1}{2}\operatorname{tr}\mathcal{M}_{\gamma(1)\leftarrow\gamma(0)}(\alpha,h),$
the eigenvalue equation can be written as
$0=\tfrac{1}{2}e^{-ik/h}\det(\mathcal{M}_{z_{0}+1\leftarrow z_{0}}(\alpha,h)-e^{ik/h})=\cos(k/h)-Z_{h}(\alpha),$ (5.62)
and we want to find $\alpha$’s for which the right hand side vanishes.
We note that
$\psi_{\gamma,\alpha}(1)-\psi_{\gamma,\alpha}(0)=c_{0}\operatorname{Im}\alpha\quad\text{where }c_{0}:=\int_{0}^{1}W(z)\,dz>0.$
Now, by Lemma 12
$\mathcal{M}_{\gamma(1)\leftarrow\gamma(0)}=B_{N}(\gamma(1);\alpha,h)^{-1}S_{\gamma}(\alpha,h)B_{N}(\gamma(0);\alpha,h)+\mathcal{O}_{\gamma,N}(h^{N+1}e^{c_{0}\operatorname{Im}\alpha/h}).$
By Lemma 14 we have
$B_{N}(\gamma(1);\alpha,h)=B_{N}(\gamma(0);\alpha,h),\quad S_{\gamma}(\alpha,h)=\begin{pmatrix}e^{-i\alpha h^{-1}c(h/\alpha)}&0\\ 0&e^{i\alpha h^{-1}c(h/\alpha)}\end{pmatrix}$.
Therefore,
$Z_{h}(\alpha)=\cos(\alpha h^{-1}c(h/\alpha))+R_{h}(\alpha),\quad R_{h}(\alpha)=\mathcal{O}_{\gamma,N}(h^{N+1}e^{c_{0}\operatorname{Im}\alpha/h}).$ (5.63)
One observation is that for the eigenvalue equation to be solvable, we need
$|\operatorname{Im}k|\leq c_{0}|\operatorname{Im}\alpha|+Ch.$ (5.64)
We can also assume that $|\operatorname{Re}k|\leq\pi h$ as conjugation by $e^{inx}$ means shifting $k$ by $nh$, $n\in\mathbb{Z}$.
We first consider the simplified equation obtained from (5.62) and (5.63):
$F_{0}(\alpha,k,h):=\cos(c(\alpha/h)\alpha/h)-\cos(k/h)=0$ (5.65)
which holds when
$c(h/\alpha)\alpha=2\pi nh\pm k,\ \ \ \ n\in\mathbb{Z},$
noting that the roots are double when $k/h\in\pi\mathbb{Z}$. The solutions satisfying (5.7) have asymptotic expansions (where we use the fact that $c_{1}=0$ in (5.35))
$\begin{split}\alpha_{\pm}(n,k,h)&=c_{0}^{-1}(2\pi nh\pm k)+d_{2}(2\pi nh\pm k)^{-1}h^{2}\\ &\ \ \ \ +d_{3}(2\pi nh\pm k)^{-2}h^{3}+\cdots,\ \ \ d_{2}=-c_{2},\cdots\end{split}$ (5.66)
We can now compare these solutions to the solutions of the equation (5.62), for $\operatorname{Im}\alpha\geq-h^{M}$, using (5.63) and Rouché’s theorem. The symmetry $\alpha\to-\alpha$ then gives the result for $\operatorname{Im}\alpha\leq 0$ as well. Also, the sign assumption is not needed as we can multiply $\alpha$ by a constant. Rescaling to eliminate $h$ we obtain the following theorem from which the rougher (no multiplicities) Theorem 3 follows:
Theorem 4.
Under the assumptions of Theorem 3 there exist a contant $C_{0}$ and maps
$\widetilde{\alpha}_{\pm}$$\mathbb{Z}\times\mathbb{C}\to\mathbb{C}$,
such that with $\alpha_{\pm}(n,k):=\alpha_{\pm}(n,k,1)$ (see (5.66)),
$\widetilde{\alpha}_{\pm}(n,k)-\alpha_{\pm}(n,k)=\mathcal{O}(|2\pi n\pm k|^{-\infty}),$
such that for $|\alpha|\geq C_{0}$, and in the notation of (5.6),
$m(\alpha,k)=|\{(n,\varepsilon):\widetilde{\alpha}_{\varepsilon}(n,k)=\alpha,\ \varepsilon=\pm 1\}|.$